### v2

# Shuchang Zhang, Hui Zhang, Hongxia Wang

Abstract
Preconditioned Proximal Point (PPP) algorithms provide a unified framework for splitting methods in image restoration. Recent advancements with RED (Regularization by Denoising) and PnP (Plug-and-Play) priors have achieved state-of-the-art performance in this domain, emphasizing the need for a meaningful particular solution. However, degenerate PPP algorithms typically exhibit weak convergence in infinite-dimensional Hilbert space, leading to uncertain solutions. To address this issue, we propose the Halpern-type Preconditioned Proximal Point (HPPP) algorithm, which leverages the strong convergence properties of Halpern iteration to achieve a particular solution. Based on the implicit regularization defined by gradient RED, we further introduce the Gradient REgularization by Denoising via HPPP called GraRED-HP3 algorithm. The HPPP algorithm is shown to have the regularity converging to a particular solution by a toy example. Additionally, experiments in image deblurring and inpainting validate the effectiveness of GraRED-HP3, showing it surpasses classical methods such as Chambolle-Pock (CP), PPP, RED, and RED-PRO.
keywords:
Halpern iteration, Preconditioned proximal point algorithms, RED, Image restoration
1 Introduction
Image restoration (IR) problems, including image deblurring, super-resolution, and inpainting, can be formulated as the following optimization problem [10, 20]:
(1) $\min_{\mathbf{x}\in\mathcal{X}}\lambda f(\mathbf{x})+g(\mathbf{K}\mathbf{x}),$
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization (or prior) to mitigate the ill-posedness of IR problems. Examples include total variation (TV) regularization $\left\|\nabla\mathbf{x}\right\|_{1}(\mathbf{K}=\nabla)$ as in [34], and sparsity regularization $\left\|\mathbf{x}\right\|_{1}(\mathbf{K}=I)$.
By the first-order optimality condition, the convex optimization problem (1) is equivalent to the following inclusion problem:
[…]
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [4, Chapter 3]. Following [20, 5], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we can reformulate (2) as:
[…]
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [1, 2]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [27]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and has been proven to converge weakly [32]. Since the resolvent $(I+\mathcal{A})^{-1}$ is generally difficult to compute, splitting methods have been proposed to address this issue. The well-known Douglas-Rachford splitting (DRS) [16] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which the resolvents $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another popular splitting algorithm is the Chambolle-Pock (CP) algorithm [10, 30], which solves the saddle-point problem of (1), i.e.,
(4) $\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}\left\langle\mathbf{K}\mathbf{x},\mathbf{y}\right\rangle+\lambda f(\mathbf{x})-g^{*}(\mathbf{y}),$
where $g^{*}:\mathcal{K}\to\mathbb{R}^{+}$ is the Fenchel conjugate functional. He and Yuan [21] were the first to analyze the Primal-Dual Hybrid Gradient (PDHG) by the preconditioned proximal point (PPP) method, given by $\mathbf{0}\in\mathcal{A}\mathbf{u}^{k+1}+\mathcal{M}(\mathbf{u}^{k+1}-\mathbf{u}^{k})$ with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Bredies et al. [6, 7, 5] developed a unified degenerate PPP algorithmic framework with a semi-definite preconditioner. Assuming that $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$ has a full domain and is single-valued, the PPP method with the classic Krasnosel’skii-Mann (KM) iteration solves for the fixed point of $\mathcal{T}$, i.e., $\mathcal{T}(\mathbf{u})=\mathbf{u}$. From the perspective of degenerate PPP [5], both DRS and CP splitting algorithms can be considered as special cases of the PPP algorithm with proper preconditioner.
However, PPP algorithms with a semi-definite preconditioner generally exhibit weak convergence in Hilbert space, resulting in uncertain solutions. It is meaningful to study the limit point of sequences generated by PPP, Bauschke et al. obtained strong convergence results of PPP in the special case of linear relation on $\mathcal{A}$ [2], the fixed-point set $\mathrm{Fix}(\mathcal{T})$ is a linear subspace, the weak limit of the PPP sequence is the $\mathcal{M}$-projection of initial point. The classic Halpern iteration [19] offers the advantage of strong convergence over other iterations (such as the KM iteration) in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [22]. This method is also known as an implicit regularization technique [15, 31]. Due to this implicit bias (projection onto solutions), degenerate PPP algorithms incorporating Halpern iteration are more likely to approximate the true solution, yielding better recovery results for IR problems. Besides the benefit of converging to a particular fixed point, Halpern iteration also possesses acceleration properties widely utilized in machine learning [15, 29]. Even without theoretical guarantees, literature [29] has demonstrated that PDHG with Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Sun et al. propose a globally convergent accelerated degenerate Proximal Point Method (dPPM) by unifying the Halpern iteration and the fast KM iteration. They further accelerate a preconditioned Alternating Direction Method of Multipliers (pADMM) via dPPM [37].
Recently, first-order optimization algorithms (steepest decent (SD), fixed-point (FP) iteration and ADMM) have played an critical role in implicit regularization, such as PnP priors [39] and RED [33], which have achieved state-of-the-art performance in IR problems [8, 44, 25, 38]. PnP methods by replacing proximal operators with denoisers do not seek the minimization of an explicit objective function, which strongly limits their interpretation [14, 25]. The RED prior defines a convex function whose gradient corresponds to the denoising residual itself, formulating a clear and well-defined objective function [14]. Thus, the gradient of RED can be used by these first-order optimization methods. To prove the global convergence of RED, based on fixed-point projection, Cohen et al. proposed the RED-PRO model [14] to bridge between RED and PnP using the hybrid steepest descent method (HSD) [42]. Assuming that the data term $f$ is a strong convexity function and the residual of the denoising neural network is nonexpansive, Ryu et al. proved the convergence of PnP-FBS(Forward-Backward Splitting) and PnP-ADMM by the classical Banach contraction principle [35], but imposing strong convexity on $f$ excludes many IR tasks [25]. To overcome the strong convexity assumption, Hurault et al. proposed proximal denoisers of PnP methods, and the gradient step denoiser [13, 24] can be interpreted as a proximal operator of implicit regularization [25]. However, these works did not obtain the particular solution. It is crucial to obtain a unique, stable, and particular solution for IR problems [17].
Since the Halpern iteration can natually obtain the particular solution (initial point projection onto the fixed-point set). Therefore, Halpern-type preconditioned proximal point called HPPP (5) is proposed to takle the above issue, i.e.,
(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}(\mathbf{u}^{k}),$
where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ is a $\mathcal{M}$-FNE operator (7) and $\{\mu_{k}\}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\mu_{k}\to 0(k\to\infty)$. In addition to TV, implicit regularization such as PnP prior is also suitable for HPPP. Under the nonexpansive assumption, we find that the residual of the denoiser can also be interpreted as the proximal operator of the implicit regularization via Moreau decomposition, making it straightforward to integrate into the primal-dual algorithm with the special semi-definte preconditioner $\mathcal{M}$. These splitting algortihms can be viewed as fixed-point iterations from the perspective of PPP. Based on HPPP and implicit regularization,  [^1] Since the gradient of RED is exactly the residual of the denoiser, we used the GraRED notion to denote the proximal operator of the implicit regularization.we propose the implicit Gradient Regularization by Denoising via HPPP called GraRED-HP3(see Algorithm 1) for IR problems. Not only can it converge to a particular solution, but it also leverages the data adaptivity of implicit regularization. This integration enriches the theoretical and algorithmic understanding of RED and PnP priors. The main contributions are as follows:
  1. 1.
    Theoretically, under the semi-definite preconditioner, we analyze the convergence of HPPP in Hilbert space. Compared with PPP, the proposed HPPP has the advantage of converging to a unique particular solution $\mathbf{u}^{*}={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(Lemma A.4), which extends the results of the classic Halpern iteration. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by HPPP; we establish an $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate for $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
  2. 2.
    Based on the special preconditioner $\mathcal{M}$, the primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}$. We apply TV or implicit regularization to HPPP, then propose the GraRED-HP3 algorithm for image restoration problems, further discuss the relationship between PnP-ADMM, GraRED-P3 (GraRED via PPP), and the proposed algorithm.
  3. 3.
    We numerically verify the regularity of HPPP with a simple 1D example, and further demonstrate the state-of-the-art performance of GraRED-HP3 on image deblurring and inpainting experiments. The results show that GraRED-HP3 outperforms the classic CP [10] and PPP [5] algorithms with TV regularization, as well as the RED [33] and RED-PRO [14] algorithms with TNRD denoiser [12].
The rest of this paper is organized as follows. In section 2, we review some useful preliminaries for convergence analysis, including the background of PPP and implicit gradient RED. In section 3, we establish the convergence and convergence rate of HPPP and apply it to TV regularization. Based on the implicit gradient RED, we propose the GraRED-HP3 algorithm for IR problems. In section 4, we verify the regularity and efficiency of the proposed algorithms with a 1D toy example. Furthermore, we validate the performance of the proposed GraRED-HP3 algorithm through image deblurring and inpainting experiments. Finally, conclusions are presented in section 5.
2 Preliminaries
[…]
2.1 Preconditioned proximal point
Bredies et al. [5] introduced a linear, bounded, self-adjoint and positive semi-definite operator admissible preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates $(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$.
Definition 2.1.
An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bouned, linear, self-adjoint, and positive semi-definite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
$\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$
is single-valued, of full-domain, and Lipschitz continuous.
Therefore, the preconditioned proximal point iteration is written into
(6) $\mathbf{u}^{0}\in\mathcal{H},\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}.$
If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly non-expansive operator [1, section 23] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is semi-definite, the property of $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [6, 9, 41, 5], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
The following notion extends monotone characteristics, i.e., $\mathcal{M}-$monotonicity.
Definition 2.2 ($\mathcal{M}$-monotonicity).
Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bouned linear positive semi-definite operator, then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0,\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
According to [5, Lemma 2.6], if $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
(7) $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|_{\mathcal{M}}^{2}+\left\|(I-\mathcal{T})\mathbf{u}-(I-\mathcal{T})\mathbf{v}\right\|^{2}_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{v}\right\|^{2}_{\mathcal{M}}.$
By the KM iteration, the relaxed PPP algorithm is written into
(8) $\mathbf{u}^{k+1}=(1-\lambda_{k})\mathbf{u}^{k}+\lambda_{k}\mathcal{T}\mathbf{u}^{k},$
where $\lambda_{k}$ satifies $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 Implicit Gradient RED
Romano et al. introduced an explicit regularizer called Regularization by Denoising (RED) [33] to construct an explicit objective function, which is defined by
(9) $g(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser ($\sigma>0$ denotes a noise level). If $D_{\sigma}(\mathbf{x})$ satisfies homogeneity $D_{\sigma}(c\mathbf{x})=cD_{\sigma}(\mathbf{x})(\forall c>0)$ and has symmetric Jacobian, then $\nabla g(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})$. Under the following nonexpansive Assumption (A[35] on residul $R=I-D_{\sigma}$, i.e.,
(A) $\left\|(\mathbf{x}-D_{\sigma}(\mathbf{x}))-(\mathbf{y}-D_{\sigma}(\mathbf{y}))\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n},$
we will prove that there exists an implicit regularization $\phi$ such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$. Here is the definition of the proximal operator and characterization of proximal operators.
Definition 2.3.
[4, Section 6] Given a function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
(10) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{x}).$
Proposition 2.4 ([28, 18]).
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c. (lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ if, and only if the following conditions hold jointly:
  1. (a)
    there exists a (convex l.s.c.) function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
  2. (b)
    $h$ is nonexpansive, i.e.,
    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
Lemma 2.5.
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, then the gradient of RED defines an implicit regularization $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ such that
(11) $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}),D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}).$
Proof 2.6.
Let $\psi(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,h(\mathbf{x})=R(\mathbf{x})$ in Proposition 2.4. By Assumption (A), there exist a function denoted $\phi^{*}$($\phi^{*}$ is the conjugate of $\phi$) such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. By Moreau decomposition [4, Theorem 6.44], $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$.
3 Halpern-type preconditioned proximal point (HPPP)
Compared to KM iteration, Halpern iteration [19] is an effective method to find a particular fixed point, i.e.,
(12) $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k},$
the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection of the initial point $\mathbf{u}^{0}$ to $\mathrm{Fix}(T)$  [19, 26, 40], i.e., $\mathbf{u}^{*}=P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$, where $P_{\Omega}(\mathbf{u}^{0})=\arg\min_{\mathbf{u}\in\Omega}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$ denotes the standard metric projection. Using the PPP to solve the inclusion problem (3), the fixed point of $\mathcal{T}$ is the solution of (3). Choosing an appropriate anchor point, such as the degraded image or an image filled with ones or zeros, can bring the projection closer to the ground-truth solution, resulting in better restoration performance. Hence, the HPPP algorithm (5) theoretically converges to the ground-truth solution under some known anchor point prior $\mathbf{a}$. When $\mathcal{T}$ satisfies the mild condition $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ [5], the HPPP algorithm is able to find a unique solution $\mathbf{u}^{*}$ of (3), where $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(see Lemma A.4).
3.1 Convergence analysis
[…]
Theorem 3.1.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
  1. (i)
    $\lim_{k\to\infty}\mu_{k}=0$;
  2. (ii)
    $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty$;
  3. (iii)
    $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0\text{ or }\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$.
Then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$ which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
Proof 3.2.
Firstly, we show that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$, where
$\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}.$
Then, we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}^{2}$
$=\mu_{k+1}^{2}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+(1-\mu_{k+1})^{2}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$
$+2\mu_{k+1}(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
$\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+\mu_{k+1}\delta_{k+1},$
where $\delta_{k}=\mu_{k}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2\mu_{k}(1-\mu_{k})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k-1}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$. By Lemma A.2, $\{\mathbf{u}^{k}\}$ is bounded, there exists $\tilde{\mathbf{u}}$ such that $\mathbf{u}^{k}\rightharpoonup\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$ according to the known condition. Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\tilde{\mathbf{u}}$ such that
$\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
$=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
Since $\mathbf{u}^{*}$ is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{a}-\mathbf{u}\right\|_{\mathcal{M}}^{2}$ according to Lemma A.4, which solves
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Therefore $\limsup_{k\to\infty}\delta_{k}\leq 0$, from Lemma A.1 we obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$. we then prove $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$.
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
it easily follows that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$ from $\mu_{k+1}\to 0$ and $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$.
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfy conditions (i)-(iii). Under mild conditions(i)-(iii), compared with [2], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point.
Corollary 3.3.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). If $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=\infty,\mu_{k}\to 0(k\to\infty),\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$, which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
Proof 3.4.
Assume that $\mathbf{u}^{k}\rightharpoonup\mathbf{u}$, $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ (Lemma A.2), it follows that
$\mathcal{T}\mathbf{u}^{k}=\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}+\mathbf{u}^{k}\rightharpoonup 0+\mathbf{u}=\mathbf{u},$
i.e., $\mathcal{T}\mathbf{u}^{k}\rightharpoonup\mathbf{u}$. From $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ and $\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}$ we have
(13) $\mathcal{A}\mathcal{T}\mathbf{u}^{k}\owns\mathcal{M}(\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k})\to 0.$
By the maximality of $\mathcal{A}$ we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ (see [1, Proposition 20.38]), hence $0\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}$ is a fixed point of $\mathcal{T}$. Thus, every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, it holds by Theorem 3.1.
Remark 3.5.
Compared Theorem 3.1, Corollary 3.3 with [5, Theorem 2.9, convergence] and [5, Corollary 2.10], $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same except we only add the mild assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$ used for convergence analysis. $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can converges weakly to the same fixed point of $\mathcal{T}$ since $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
3.2 Convergence rate
In this section, we will give the result of the sub-linear convergence rate of the gap between two successive iterations. We now present a technical Lemma [36] for convergence rate.
Lemma 3.6 ([36]).
Let $M>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers which satisfy $a_{1}<M$ and
$a_{k+1}\leq(1-\gamma b_{k+1})a_{k}+(b_{k}-b_{k+1})c_{k},k\geq 1,$
where $\gamma\in(0,1]$, $\{b_{k}\}_{k\in\mathbb{N}}$ is a sequence which is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M<\infty$. Then, the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
$a_{k}\leq\frac{MJ}{\gamma k},k\geq 1,$
where $J=\lfloor\frac{2}{\gamma}\rfloor$.
Corollary 3.7.
Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
  1. (i)
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2M^{\prime}}{k}$, i.e., $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$;
  2. (ii)
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
Proof 3.8.
For (i), denote $M^{\prime}=\sup_{k\in\mathbb{N}}\{\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}$, since
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M^{\prime},$
it derives (i) by Lemma 3.6. For (ii), since
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|$
$\leq\frac{2M^{\prime}C}{k-1}+\frac{2M^{\prime}}{k(k+1)},k\geq 2.$
and thus $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
3.3 HPPP with TV regularization
Chambolle-Pock(CP) is a first-order primal-dual algorithm to solve non-smooth convex optimization problems with known saddle-point structure [10]. Take the starting point $(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$, the CP primal-dual algorithm is written into the following form:
(14) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=(I+s\partial g^{*})^{-1}\left(\mathbf{y}_{k}+s\mathbf{K}(2\mathbf{x}_{k+1}-\mathbf{x}_{k})\right),\end{array}\right.$
here we consider TV regularization $\mathbf{K}=\nabla$. The convergence of CP algorithm was proved when step sizes satisfy $\tau s\left\|\mathbf{K}\right\|^{2}<1$. By introducing the operator $\mathcal{A}$ and preconditioner $\mathcal{M}$, set
(15) $\mathcal{A}=\begin{pmatrix}\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}.$
As fistly noticed in [20, 5], the CP method is a special PPP method $0\in\mathcal{A}\mathbf{x}^{k+1}+\mathcal{M}(\mathbf{x}^{k+1}-\mathbf{x}^{k})$. Bredies et al. proved the weak convergence of the CP algorithm about the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ [5], i.e., $\mathcal{M}$ is a semi-definite preconditioner. To apply HPPP with TV regularization, under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ we obtain
(16) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\frac{2s\mathbf{K}(\mathbf{x}_{k+1}-\mu_{k+1}\mathbf{x}_{a})}{1-\mu_{k+1}}-s\mathbf{K}\mathbf{x}_{k}+\mathbf{y}_{k}\right),\end{array}\right.$
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points, respectively, and $\mu_{k}$ satisfies conditions(i)-(iii) of Theorem 3.1.
3.4 HPPP with implicit regularization
Recently, RED frameworks have achieved state-of-the-art performance for inverse problems by utilizing CNN (convolutional neural network) denoiser as regularization [14], such as DnCNN [46]. Ryu et al. proposed a nonexpansive residual of DnCNN to ensure the convergence of PnP methods [35]. Gradient step denoisers, defined as $D_{\sigma}=I-\nabla g_{\sigma}$ for Gaussian noise level $\sigma$, have been proposed to generalize RED [13, 24], where $g_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}$ is a scalar function parameterized by a differentiable neural network. The gradient step denoiser can be interpreted as the proximal operator of an implicit regularization [25].
The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$ to solve (3). According to subsection 2.2, $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, we consider the special application $g=\phi,\mathbf{K}=I$ with the degenerate $\mathcal{M}(\tau s=1)$.
The proposed HPPP algorithm has the advantage of converging to a particular solution than PPP. Due to the data adaptability of CNN, which has the powerful expressiveness to represent the implicit regularization $\phi$ [44, 45]. Based on HPPP and GraRED, we propose the implicit Gradient RED via HPPP called GraRED-HP3 (see Algorithm 1). According to (8), we further discuss GraRED-P3 Algorithm 2 for IR problems.
PnP-ADMM [39, 11, 35] is a well-known optimization method to solve (1) when $g=\phi$, i.e.,
(17) $\mathbf{z}^{k+1}=$ $\mathrm{prox}_{\phi}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k})=D_{\sigma}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k}),$
$\mathbf{v}^{k+1}=$ $\mathrm{prox}_{\lambda f}(\mathbf{z}^{k+1}+\bar{\mathbf{u}}^{k}),$
$\bar{\mathbf{u}}^{k+1}=$ $\bar{\mathbf{u}}^{k}+\mathbf{z}^{k+1}-\mathbf{v}^{k+1}.$
Let $\mathbf{w}^{k}=\mathbf{v}^{k}+\bar{\mathbf{u}}^{k}$, then PnP-ADMM can be written into equivalent Douglas-Rachford Splitting (DRS) form:
(18) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k}).$
Using $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\lambda_{k}=1$ in GraRED-P3 results exactly in the above DRS iteration. Thus, PnP-ADMM can be obtained from the perspective of PPP [6, 5], and its convergence follows by [5, Corollary 2.15].
The following Theorem 3.9 verifies the convergence of Algorithm 2 and Algorithm 2.
Theorem 3.9.
Assume that a denoiser $D_{\sigma}:\mathcal{H}\to\mathcal{H}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, relaxing parameter $20<\inf_{k}\lambda_{k}\leq\sup_{k}\lambda_{k}<$ 2 and RED is convex, then $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 2 converges to some uncertain solution $\mathbf{x}^{*}$ of (1), while $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 converges to the particular solution $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of (3).
Proof 3.10.
Followed by [5, Corollary 2.10] and Corollary 3.3.
Algorithm 1 GraRED-HP3
1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, diminishing anchor coefficient $\mu_{k}\in[0,1],\mu_{k}\to 0(k\to\infty),\sum_{k=0}^{\infty}\mu_{k}=+\infty$, and $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.
2:for $k=0,1,2,\cdots,N-1$ do
3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
4:     $\mathbf{x}^{k+1}=\mu_{k}\mathbf{x}_{a}+(1-\mu_{k})\mathbf{d}^{k}$
5:     $\mathbf{v}^{k+1}=R\left(s(2\mathbf{d}^{k}-\mathbf{x}^{k})+\mathbf{y}^{k}\right)$
6:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})\mathbf{v}^{k+1}$
7:end for
8:$\mathbf{x}^{N}$.
Algorithm 2 GraRED-P3
1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, $\mathcal{A}=\begin{pmatrix}\lambda\partial f&{I}\\ -I&\partial{\phi}^{*}\end{pmatrix}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.
2:for $k=0,1,2,\cdots,N-1$ do
3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
4:     $\mathbf{x}^{k+1}=\lambda_{k}\mathbf{d}^{k}+(1-\lambda_{k})\mathbf{x}^{k}$
5:     $\mathbf{y}^{k+1}=\lambda_{k}R(\mathbf{y}^{k}+s(2\mathbf{d}^{k}-\mathbf{x}^{k}))+(1-\lambda_{k})\mathbf{y}^{k}$
6:end for
7:$\mathbf{x}^{N}$.
4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. Firstly, we will verify the regularity of HPPP by an easy 1D example. Then, we compare the CP (14), PPP (8), GraRED-HP3 algorithms (see Algorithm 1), and GraRED-P3 (see Algorithm 2) for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms.
4.1 A toy example
We consider the following optimization problem in $\mathbb{R}$, i.e.,
(19) $\min_{x\in\mathbb{R}}f(x)+g(x),$
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$ (See Figure 1).
Figure 1: Plot $f(x)+g(x)$.
Since for any $y\in\mathbb{R}$,
$\min\{(1+y)x-1,yx\}=\begin{cases}(1+y)x-1,&x<1\\ yx,&x\geq 1\end{cases},$
then $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y),y\in\mathbb{R}$, where
$\delta_{C}(y)=\begin{cases}0,&y\in C,\\ \infty,&y\notin C.\end{cases}$
Therefore, the corresponding saddle-point problem is:
$\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}xy+f(x)-g^{*}(y).$
We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fixed $x^{*}\geq 1$, then
$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0;\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$ and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
For this toy example shown in Figure 2, we choose the same initial point $\mathbf{u}^{0}=(-6,6)$ or $(0,0)$, three different anchor points $\mathbf{a}=(12,10),(12,9),(12,8)$, the semi-definite $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, and the total itertation number $N=1000$. As shown in Figure 2, $\mathbf{u}^{k}$ generated by HPPP (5) can converge to the particular saddle point controlled by the anchor point. While the limit of $\mathbf{u}^{k}$ generated by PPP (8) may be uncertain, which is related with $\mathbf{u}^{0}$ and $\lambda_{k}$. For example, the sequence $\mathbf{u}^{k}$ oscillates around the limit $(1.8,0)$ shown in Figure 2(d).
Figure 2: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ in the $xOy$ Cartesian coordinate system.
As shown in Figure 3, $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}=\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be geometrically interpreted as $\mathcal{M}$-projection, i.e., the point $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ projection onto the line $x-y=x_{a}-y_{a}$ through the anchor point $\mathbf{a}=(x_{a},y_{a})$. Therefore, the $\mathcal{M}$-projection of $\mathbf{a}=(x_{a},y_{a})$ onto $\Omega$ is definitely solved by
$P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0;\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
Figure 3: $\mathcal{M}$-projection onto the saddle point set.
4.2 Image deblurring
Firstly, we compared with classic TV regularization about CP [10], PPP [5] , and HPPP(see details in 3.3). For classic TV -$\ell^{2}$ regularization,
(20) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|_{2}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator. In case $\mathbf{A}\mathbf{x}$ can be written as a convolution, i.e., $\mathbf{A}\mathbf{x}=\mathbf{k}\ast\mathbf{x}$, where $\mathbf{k}$ is the blurring convolution kernel. The primal-dual model of TV -$\ell^{2}$ deblurring is
(21) $\min_{\mathbf{u}\in\mathbb{R}^{n}}\max_{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{u},\mathrm{div}\mathbf{p}\rangle+\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2}-\delta_{P}(\mathbf{p}),$
where $f(\mathbf{u})=\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm defined as:
$\left\|\mathbf{p}\right\|_{\infty}=\max_{i,j}|\mathbf{p}_{i,j}|,|\mathbf{p}_{i,j}|=\sqrt{(\mathbf{p}_{i,j}^{1})^{2}+(\mathbf{p}_{i,j}^{2})^{2}}.$
We can calculate the proximal operator $(I+\tau\partial g^{*})^{-1}$ of the indicator $g^{*}(\mathbf{p})$, i.e.,
$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}.$
The resolvent operator for $f(\mathbf{x})$ can be computed by FFT.
$\mathbf{x}$ $=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})$
$=\arg\min_{\mathbf{x}}\frac{\|\mathbf{x}-\tilde{\mathbf{x}}\|}{2\tau}+\frac{\lambda}{2}\left\|\mathbf{k}*\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$
$=\mathcal{F}^{-1}\left(\frac{\tau\lambda\mathcal{F}(\mathbf{y})\mathcal{F}^{*}\left(\mathbf{k}\right)+\mathcal{F}(\tilde{\mathbf{u}})}{\tau\lambda\mathcal{F}\left(\mathbf{k}\right)^{2}+1}\right),$
where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the FFT and inverse FFT, respectively.
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive WGN with noise level $0.01$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $\left\|K\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
  • GraRED-P3 : $\tau=1,s=1,\lambda_{k}=0.2,\lambda=20$;
  • GraRED-HP3 : $\tau=1,s=1,\mu_{k}=1/(k+2),\lambda=20,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
Secondly, we compared with RED and RED-PRO. The RED-PRO model uses the hybrid steepest descent method (HSD) [42, 43] to solve IR problems. Following [33, 14], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with an additive WGN with noise level $\sigma=\sqrt{2}$. The original RGB image is converted to the YCbCr image, PnP restoration algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images.
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3. The performance of three different methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularization is more powerful to regularize inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 4. To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 5, we plot their respective evolutions of PSNR values for iterations for the image House with 10 random initial points. The HPPP algorithm converges faster (less than 200 iterations) and achieves better PSNR values than CP and PPP algorithms. Once the anchor point is chosen, the proposed algorithm is more robust than CP and PPP algorithms for image deblurring with random initializations. To verify Corollary 3.7, we show the trend of convergence rate of the gap $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ between two successive iterations throughout the iterations with step size $\min\{\frac{2}{k},1\}(k\geq 1)$ in Figure 6.
Table 1: Parameters on CP, PPP, and HPP with TV regularization for image deblurring.
CP $\tau=s=1/K=0.57$
PPP $\tau=s=1/K=0.57,\lambda_{k}=1.95$ or $\lambda_{k}=1.2$
HPPP $\tau=s=1/K=0.57,\mu_{k}=\frac{1}{k+2},\mathbf{x}_{a}=\mathbf{A}^{\mathrm{T}}\mathbf{y},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$
Table 2: Deblurring results of gray images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
Cameraman House Pepper Starfish Butterfly Craft Parrots Barbara Boat
CP 26.04 30.86 25.99 27.53 27.85 25.51 26.88 24.46 28.99
PPP 26.05 30.85 25.99 27.53 27.85 25.51 26.88 24.46 28.98
HPPP(TV) 26.00 31.39 26.05 27.65 27.99 25.51 26.82 24.51 29.09
GraRED-P3 26.87 32.51 28.34 29.16 29.70 26.90 27.95 24.66 29.68
GraRED-HP3 26.85 32.51 28.80 29.15 29.67 27.03 27.89 24.66 29.99
Refer to captionRefer to caption
(a) Clean image
Refer to captionRefer to caption
(b) Degraded image
Refer to captionRefer to caption
(c) CP ($27.85$dB)
Refer to captionRefer to caption
(d) $\lambda_{k}=1.95$: PPP ($27.85$dB)
Refer to captionRefer to caption
(e) $\lambda_{k}=1.2$: PPP ($27.85$dB)
Refer to captionRefer to caption
(f) HPPP ($27.65$dB)
Refer to captionRefer to caption
(g) GraRED-P3 ($29.70$dB)
Refer to captionRefer to caption
(h) GraRED-HP3 ($29.67$dB)
Figure 4: Deblurring result of Butterfly degraded with the Gaussian blur kernel with noise level $\sigma=0.01$.
Figure 5: Random initialization.
Figure 6: The convergence rate of $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ of GraRED-H3. The image Starfish is degraded by a Gaussian PSF with noise level $0.01$.
From the deblurring experiment Table 3 and Figure 7, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective.
Table 3: Recovery results are obtained by RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3 with TNRD [12].
Algorithms Uniform kernel Gaussian kernel($\sigma_{k}=1.6$)
Bike Butterfly Flower Hat Bike Butterfly Flower Hat
RED [33] 26.10 30.41 30.18 32.16 27.90 31.66 32.05 33.30
RED-PRO(HSD) [14] 24.95 27.24 29.38 31.55 27.36 30.55 31.81 33.07
RRP [14] 26.48 30.64 30.46 32.25 28.02 31.66 32.08 33.26
GraRED-P3 26.55 30.72 30.67 32.43 28.13 31.81 32.42 33.51
GraRED-HP3 26.80 30.88 30.74 32.42 28.06 31.80 32.27 33.50
Figure 7: Deblurring result of Butterfly degraded by uniform kernel.
4.3 Image inpainting
In this section, we use the proposed algorithm to solve TV image inpainting problems and compare their numerical results with CP [10], PPP [5], and HPPP algorithms. The discrete image inpainting model is
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
where $\left\|\nabla\mathbf{x}\right\|_{1}$ is the TV regularization, $\lambda,\beta$ are balance coefficients, $\odot$ indicates pointwise multiplication. Then the saddle-point problem is
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\max_{p\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{x},\mathrm{div}\mathbf{p}\rangle+\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}-\delta_{P}(\mathbf{p}),$
where $f(\mathbf{x})=\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm. Their resolvent operators of $f,g^{*}$ are
$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}$
and
$\mathbf{x}=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})=\frac{2\tau\mathbf{M}\odot\mathbf{y}+\tilde{\mathbf{x}}}{1+2\tau\mathbf{M}},$
the multiplication and division operators should be understood pointwise in the above formula.
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start their iterations with the degraded images. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters :
  • HPPP: $\tau=s=1/\left\|K\right\|=0.57$, anchor point $\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$, stepsize $\mu_{k}=\frac{1}{10(k+2)}$;
  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$ or $\lambda_{k}=1.2$;
  • CP: $\tau=s=1/\left\|K\right\|=0.57$.
Both GraRED-P3 and GraRED-HP3 use DnCNN [35], other parameters are used below:
  • GraRED-P3 : $\tau=10,s=0.1,\lambda_{k}=0.2,\lambda=5$;
  • GraRED-HP3 : $\tau=10,s=0.1,\mu_{k}=0.05/(k+2),\lambda=5,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
In Table 4 and Table 5, we compared the numerical performance of classic algorithms with TV regularization. As we can see from both two tables, the proposed two algorithms outperform other algorithms. In Figure 8 and 9, we compare visualization results of House degraded by Bernoulli random mask and character mask, the proposed algorithms achieve better visual performance than TV regularization.
Moreover, we compare the recovery results about different anchors $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})$ with fixed $\mathbf{y}_{a}=\mathbf{0}$ and step size $\frac{1}{k+1}$. As shown in Figure 10, the anchor $\mathbf{x}_{a}=\mathbf{1}$ can achieve the best peroformance for random inpainting, which illustates the projection point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of $\mathbf{a}=(\mathbf{1},\mathbf{0})$ is closest to the true solution. Anchors selection is not difficult, GraRED-HP3 can achieve similar peroformance with other anchors.
Table 4: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, Bernoulli random missing).
Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
CP 23.54 28.94 24.53 23.89 23.59 23.42 23.20 23.10 26.29
PPP 23.58 29.05 24.53 24.32 23.76 23.43 23.33 23.13 26.36
HPPP(TV) 23.89 29.19 24.55 24.46 24.04 23.52 23.44 23.35 26.45
GraRED-P3 29.92 36.15 32.67 31.99 31.95 29.53 30.05 32.25 32.91
GraRED-HP3 29.90 36.14 32.68 31.97 31.92 29.52 30.04 32.23 32.90
Table 5: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, character texture missing).
Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
CP 25.87 31.58 29.40 26.16 25.60 26.33 25.20 24.21 25.23
PPP 26.33 31.78 30.13 26.86 26.08 26.47 25.64 27.32 28.05
HPPP 26.33 32.00 30.02 26.88 26.38 26.51 25.83 27.34 28.09
GraRED-P3 29.18 37.46 33.79 30.89 29.45 29.20 26.04 29.68 30.04
GraRED-HP3 29.13 37.47 33.80 30.93 29.42 29.20 26.02 29.60 30.04
Refer to captionRefer to caption
(a) Clean image
Refer to captionRefer to caption
(b) Degraded image
Refer to captionRefer to caption
(c) CP ($28.94$dB)
Refer to captionRefer to caption
(d) $\lambda_{k}=1.6$: PPP ($29.06$dB)
Refer to captionRefer to caption
(e) $\lambda_{k}=1.2$: PPP ($29.09$dB)
Refer to captionRefer to caption
(f) HPPP ($29.20$dB)
Refer to captionRefer to caption
(g) GraRED-P3 ($36.15$dB)
Refer to captionRefer to caption
(h) GraRED-HP3 ($36.14$dB)
Figure 8: Recovery results of House degraded by random mask.
Refer to captionRefer to caption
(a) Clean image
Refer to captionRefer to caption
(b) Degraded image
Refer to captionRefer to caption
(c) CP ($31.58$dB)
Refer to captionRefer to caption
(d) $\lambda_{k}=1.6$: PPP ($31.79$dB)
Refer to captionRefer to caption
(e) $\lambda_{k}=1.2$: PPP ($31.74$dB)
Refer to captionRefer to caption
(f) HPPP ($32.00$dB)
Refer to captionRefer to caption
(g) GraRED-P3 ($37.46$dB)
Refer to captionRefer to caption
(h) GraRED-HP3 ($37.47$dB)
Figure 9: Recovery results of House degraded by character mask.
(a) Clean image
(b) Degraded image
(c) Anchor 1: 31.89dB
(d) Anchor 2: 31.95dB
(e) Anchor 3: 31.97dB
(f) Anchor 4: 31.85dB
Figure 10: Compared with different anchors for random inpainting, i.e., $\mathbf{x}_{a}=\mathbf{0},\mathbf{M}\odot\mathbf{x}_{0}+0.5(1-\mathbf{M}),\mathbf{1},\mathbf{x}_{0}$.
5 Conclusions
In this paper, based on Halpern iteration and gradient RED, we propose HPPP and GraRED-HP3 for IR problems, which can converge strongly to a particular fixed point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$. Numerical experiments verify the regularity of HPPP, and the effectiveness GraRED-HP3 for image deblurring and inpainting, which can achieve better performance than classic algorithms with TV regularization. In the future, we plan to study the convergence of nonconvex implicit regularization [25] and extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [3].
Appendix A Boundedness and $\mathcal{M}$-projection
Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important Lemma from [23, Lemma 2.5].
Lemma A.1.
Let $\{a_{k}\}_{k\in\mathbb{N}}$be a sequence of non-negative real numbers satisfying
(22) $a_{k+1}\leq(1-\mu_{k})a_{k}+\mu_{k}\beta_{k}+\gamma_{k},$
where $\{\mu_{k}\}_{k\in\mathbb{N}},\{\beta_{k}\}_{k\in\mathbb{N}},\{\gamma_{k}\}_{k\in\mathbb{N}}$ satisfies the following conditions:
  1. (i)
    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivently $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
  2. (ii)
    $\lim\sup_{k\to\infty}\beta_{k}\leq 0$;
  3. (iii)
    $\gamma_{k}\geq 0,\sum_{k=0}^{\infty}\gamma_{k}<\infty$.
Then $\lim_{k\to\infty}a_{k}=0$.
Lemma A.2.
Let $\mathcal{T}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{\mathbf{u}^{k}\}$ be the sequence generated by (5), $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$, and $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, which satisfies the following:
  1. (i)
    $\{\mathbf{u}^{k}\},\{\mathcal{T}\mathbf{u}^{k}\}(k\in\mathbb{N})$ are bounded;
  2. (ii)
    $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0,k\to\infty$;
  3. (iii)
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0,\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0,k\to\infty$.
Proof A.3.
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$, $C=L\left\|\mathcal{C}\right\|$. For $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$,
$\left\|\mathcal{T}\mathbf{u}^{\prime}-\mathcal{T}\mathbf{u}^{\prime\prime}\right\|$ $=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime}-(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime\prime}\right\|$
$\leq L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
For (i), $\mathbf{u}^{k}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$,
$\left\lVert\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\rVert_{\mathcal{M}}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}$
$\leq\dots\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}.$
Furthermore,
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\cdot C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}<+\infty.$
and
$\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}<+\infty$
So is the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
For (ii), we should show $\mathcal{T}$ is $\mathcal{M}$-asymptotically regular $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$.
$\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$. There exists a positive real number $M>0$ such that
$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq M,$
it follows from the $\mathcal{M}$-FNE operator $\mathcal{T}$ and boundedness of $\{\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}_{k\in\mathbb{N}}$,
(23) $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
$\leq(1-\mu_{k+1})\left\|\mathcal{T}u^{k}-\mathcal{T}u^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$
$\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M$
If $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$ from Lemma A.1.
(24) $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}+\mathbf{u}^{k+1}-\mathcal{T}u^{k}\right\|_{\mathcal{M}}$
$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
Thus, $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$.
For (iii), there exists $M^{\prime}$ such that $\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|\leq M^{\prime}$,
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
$\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
Therefore $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0$.
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\to 0,k\to\infty.$
Lemma A.4.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$, and $\mathrm{Fix}(\mathcal{T})$ be a closed convex subset of a Hilbert space $\mathcal{H}$, and $\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function. There exists a unique solution $u^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
(25) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Proof A.5.
Let $l(\mathbf{u})=\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is the differentiable convex function. Assume that $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ is the optimal solution, $\mathrm{Fix}(\mathcal{T})$ is the convex set, thus $t\in(0,1),\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for $\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$,
$\lim_{t\to 0}\frac{l(\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*}))-l(\mathbf{u}^{*})}{t}$ $=\left\langle l^{\prime}(\mathbf{u}^{*}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
$=\left\langle 2\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
$=2\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
If $\mathbf{u}^{**}$ is the another solution such that $\left\langle\mathbf{u}^{**}-a,\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0$. Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in the above two inequalities, respectively.
(26) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
(27) $\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
If we add two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. Since
$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$
It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$ from above inequality.
As mentioned in [2, 9] and uniqueness projection onto $\mathrm{Fix}(\mathcal{T})$, we can introduce the following notion of $\mathcal{M}$-projection.
Definition A.6 ($\mathcal{M}$-projection).
Assume $\forall\mathbf{u}^{0}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{u}^{0}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{u}^{0}\right\|_{\mathcal{M}}(\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}))$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}-$projection of $\mathbf{u}^{0}$ onto $\mathrm{Fix}(\mathcal{T})$, also denoted $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$.
Lemma A.7.
Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ and $\mathbf{u}^{0}\in\mathcal{H}$, then the following conditions are equivalent:
  1. (i)
    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$;
  2. (ii)
    $\left\langle\mathbf{u}^{*}-\mathbf{u}^{0},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
Proof A.8.
See Lemma A.4 and Definition A.6.

### v4

@@
 Abstract
-Preconditioned Proximal Point (PPP) algorithms provide a unified framework for splitting methods in image restoration.
-Recent advancements with RED (Regularization by Denoising) and PnP (Plug-and-Play) priors have achieved state-of-the-art performance in this domain, emphasizing the need for a meaningful particular solution.
-However, degenerate PPP algorithms typically exhibit weak convergence in infinite-dimensional Hilbert space, leading to uncertain solutions.
-To address this issue, we propose the Halpern-type Preconditioned Proximal Point (HPPP) algorithm, which leverages the strong convergence properties of Halpern iteration to achieve a particular solution.
-Based on the implicit regularization defined by gradient RED, we further introduce the Gradient REgularization by Denoising via HPPP called GraRED-HP3 algorithm.
-The HPPP algorithm is shown to have the regularity converging to a particular solution by a toy example.
-Additionally, experiments in image deblurring and inpainting validate the effectiveness of GraRED-HP3, showing it surpasses classical methods such as Chambolle-Pock (CP), PPP, RED, and RED-PRO.
+Recently, the degenerate preconditioned proximal point (PPP) method provides a unified and flexible framework for designing and analyzing operator-splitting algorithms such as Douglas-Rachford (DR).
+However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants.
+To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method.
+Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors such as Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method.
+Finally, numerical experiments including several toy examples and image restoration validate the effectiveness of our proposed algorithms.
 keywords:
-Halpern iteration, Preconditioned proximal point algorithms, RED, Image restoration
+Halpern iteration, Preconditioned proximal point algorithms, Plug-and-Play Prior, Regularization by Denoising, Image restoration
 1 Introduction
@@
 Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces.
-The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization (or prior) to mitigate the ill-posedness of IR problems.
-Examples include total variation (TV) regularization $\left\|\nabla\mathbf{x}\right\|_{1}(\mathbf{K}=\nabla)$ as in [34], and sparsity regularization $\left\|\mathbf{x}\right\|_{1}(\mathbf{K}=I)$.
+The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization such as TV [48] to mitigate the ill-posedness of IR problems.
 By the first-order optimality condition, the convex optimization problem (1) is equivalent to the following inclusion problem:
 […]
-where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [4, Chapter 3].
-Following [20, 5], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we can reformulate (2) as:
+where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3].
+Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as:
 […]
@@
 When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39].
-The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and has been proven to converge weakly [32].
-Since the resolvent $(I+\mathcal{A})^{-1}$ is generally difficult to compute, splitting methods have been proposed to address this issue.
-The well-known Douglas-Rachford splitting (DRS) [16] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which the resolvents $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain.
-Another popular splitting algorithm is the Chambolle-Pock (CP) algorithm [10, 30], which solves the saddle-point problem of (1), i.e.,
+The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46].
+However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue.
+The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain.
+Another way to solve (1) is to transform it into the saddle-point problem [16, 43], i.e.,
 (4) $\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}\left\langle\mathbf{K}\mathbf{x},\mathbf{y}\right\rangle+\lambda f(\mathbf{x})-g^{*}(\mathbf{y}),$
-where $g^{*}:\mathcal{K}\to\mathbb{R}^{+}$ is the Fenchel conjugate functional.
-He and Yuan [21] were the first to analyze the Primal-Dual Hybrid Gradient (PDHG) by the preconditioned proximal point (PPP) method, given by $\mathbf{0}\in\mathcal{A}\mathbf{u}^{k+1}+\mathcal{M}(\mathbf{u}^{k+1}-\mathbf{u}^{k})$ with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
-Bredies et al.
-[6, 7, 5] developed a unified degenerate PPP algorithmic framework with a semi-definite preconditioner.
-Assuming that $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$ has a full domain and is single-valued, the PPP method with the classic Krasnosel’skii-Mann (KM) iteration solves for the fixed point of $\mathcal{T}$, i.e., $\mathcal{T}(\mathbf{u})=\mathbf{u}$.
-From the perspective of degenerate PPP [5], both DRS and CP splitting algorithms can be considered as special cases of the PPP algorithm with proper preconditioner.
-However, PPP algorithms with a semi-definite preconditioner generally exhibit weak convergence in Hilbert space, resulting in uncertain solutions.
-It is meaningful to study the limit point of sequences generated by PPP, Bauschke et al.
-obtained strong convergence results of PPP in the special case of linear relation on $\mathcal{A}$ [2], the fixed-point set $\mathrm{Fix}(\mathcal{T})$ is a linear subspace, the weak limit of the PPP sequence is the $\mathcal{M}$-projection of initial point.
-The classic Halpern iteration [19] offers the advantage of strong convergence over other iterations (such as the KM iteration) in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [22].
-This method is also known as an implicit regularization technique [15, 31].
-Due to this implicit bias (projection onto solutions), degenerate PPP algorithms incorporating Halpern iteration are more likely to approximate the true solution, yielding better recovery results for IR problems.
-Besides the benefit of converging to a particular fixed point, Halpern iteration also possesses acceleration properties widely utilized in machine learning [15, 29].
-Even without theoretical guarantees, literature [29] has demonstrated that PDHG with Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction.
-Sun et al.
-propose a globally convergent accelerated degenerate Proximal Point Method (dPPM) by unifying the Halpern iteration and the fast KM iteration.
-They further accelerate a preconditioned Alternating Direction Method of Multipliers (pADMM) via dPPM [37].
-Recently, first-order optimization algorithms (steepest decent (SD), fixed-point (FP) iteration and ADMM) have played an critical role in implicit regularization, such as PnP priors [39] and RED [33], which have achieved state-of-the-art performance in IR problems [8, 44, 25, 38].
-PnP methods by replacing proximal operators with denoisers do not seek the minimization of an explicit objective function, which strongly limits their interpretation [14, 25].
-The RED prior defines a convex function whose gradient corresponds to the denoising residual itself, formulating a clear and well-defined objective function [14].
-Thus, the gradient of RED can be used by these first-order optimization methods.
-To prove the global convergence of RED, based on fixed-point projection, Cohen et al.
-proposed the RED-PRO model [14] to bridge between RED and PnP using the hybrid steepest descent method (HSD) [42].
-Assuming that the data term $f$ is a strong convexity function and the residual of the denoising neural network is nonexpansive, Ryu et al.
-proved the convergence of PnP-FBS(Forward-Backward Splitting) and PnP-ADMM by the classical Banach contraction principle [35], but imposing strong convexity on $f$ excludes many IR tasks [25].
-To overcome the strong convexity assumption, Hurault et al.
-proposed proximal denoisers of PnP methods, and the gradient step denoiser [13, 24] can be interpreted as a proximal operator of implicit regularization [25].
-However, these works did not obtain the particular solution.
-It is crucial to obtain a unique, stable, and particular solution for IR problems [17].
-Since the Halpern iteration can natually obtain the particular solution (initial point projection onto the fixed-point set).
-Therefore, Halpern-type preconditioned proximal point called HPPP (5) is proposed to takle the above issue, i.e.,
-(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}(\mathbf{u}^{k}),$
-where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ is a $\mathcal{M}$-FNE operator (7) and $\{\mu_{k}\}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\mu_{k}\to 0(k\to\infty)$.
-In addition to TV, implicit regularization such as PnP prior is also suitable for HPPP.
-Under the nonexpansive assumption, we find that the residual of the denoiser can also be interpreted as the proximal operator of the implicit regularization via Moreau decomposition, making it straightforward to integrate into the primal-dual algorithm with the special semi-definte preconditioner $\mathcal{M}$.
-These splitting algortihms can be viewed as fixed-point iterations from the perspective of PPP.
-Based on HPPP and implicit regularization,  [^1] Since the gradient of RED is exactly the residual of the denoiser, we used the GraRED notion to denote the proximal operator of the implicit regularization.we propose the implicit Gradient Regularization by Denoising via HPPP called GraRED-HP3(see Algorithm 1) for IR problems.
-Not only can it converge to a particular solution, but it also leverages the data adaptivity of implicit regularization.
-This integration enriches the theoretical and algorithmic understanding of RED and PnP priors.
+where $g^{*}:\mathcal{K}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$.
+In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) splitting method [16, 43] and the Primal-Dual Hybrid Gradient (PDHG) method [27], are designed and studied extensively.
+In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
+Recently, Bredies et al.
+[11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semi-definite $\mathcal{M}$.
+By choosing appropriate preconditioners, this framework could cover the DR and CP splitting algorithms [9].
+It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34].
+For example, Bauschke et al.
+demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3].
+In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point.
+Moreover, Sun et al.
+[52] proposed an accelerated preconditioned Alternating Direction Method of Multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6].
+Similarly, Chen et al.
+introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semi-proximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38].
+A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8].
+Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces.
+Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
+The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29].
+Then the Halpern iteration is also known as an implicitly regularized method [22].
+Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems.
+Achieving stable reconstruction for ill-posed inverse problems is important [24].
+Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29].
+Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction.
+Based on the Halpern iteration, the Halpern-type preconditioned proximal point algorithm called HPPP (5) is thus proposed to overcome the two limitations of the degenerate PPP method,
+(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k},$
+where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
+PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53].
+Buzzard et al.
+provided a Consensus Equilibrium interpretation on denoiser priors [14].
+Romano et al.
+introduced RED (Regularization by Denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization.
+Powerful denoisers (e.g., DnCNN [66]) typically do not meet the conditions of RED [45].
+To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45].
+Based on the fixed-point projection, Cohen et al.
+proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model.
+Meanwhile, Ryu et al.
+proved the convergence of PnP-FBS and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49].
+However, enforcing strong convexity on $f$ precludes many IR tasks [33].
+Furthermore, Cohen et al.
+parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault et al.
+proposed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33].
+Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on their equivalence [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
+Based on HPPP and denoiser priors,  [^1] Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.we propose the Gradient Regularization by Denoising via HPPP called GraRED-HP3 for IR problems.
 The main contributions are as follows:
   1. 1.
-    Theoretically, under the semi-definite preconditioner, we analyze the convergence of HPPP in Hilbert space.
-Compared with PPP, the proposed HPPP has the advantage of converging to a unique particular solution $\mathbf{u}^{*}={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(Lemma A.4), which extends the results of the classic Halpern iteration.
-Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by HPPP; we establish an $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate for $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
+    Theoretical Contributions.
+The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semi-definite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution.
+Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
   2. 2.
-    Based on the special preconditioner $\mathcal{M}$, the primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}$.
-We apply TV or implicit regularization to HPPP, then propose the GraRED-HP3 algorithm for image restoration problems, further discuss the relationship between PnP-ADMM, GraRED-P3 (GraRED via PPP), and the proposed algorithm.
+    Algorithmic Development.
+We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint.
+Noting that the PnP-ADMM algorithm can be reformulated as an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
   3. 3.
-    We numerically verify the regularity of HPPP with a simple 1D example, and further demonstrate the state-of-the-art performance of GraRED-HP3 on image deblurring and inpainting experiments.
-The results show that GraRED-HP3 outperforms the classic CP [10] and PPP [5] algorithms with TV regularization, as well as the RED [33] and RED-PRO [14] algorithms with TNRD denoiser [12].
+    Experimental Validation.
+We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency with several toy examples.
+In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
 The rest of this paper is organized as follows.
-In section 2, we review some useful preliminaries for convergence analysis, including the background of PPP and implicit gradient RED.
-In section 3, we establish the convergence and convergence rate of HPPP and apply it to TV regularization.
-Based on the implicit gradient RED, we propose the GraRED-HP3 algorithm for IR problems.
-In section 4, we verify the regularity and efficiency of the proposed algorithms with a 1D toy example.
-Furthermore, we validate the performance of the proposed GraRED-HP3 algorithm through image deblurring and inpainting experiments.
+In section 2, we review some preliminaries for convergence analysis.
+In section 3, we analyze the convergence of HPPP.
+Combining HPPP and denoiser priors, we propose GraRED-HP3 for IR problems.
+In section 4, we verify the advantages of HPPP with several toy examples.
+Furthermore, we validate the performance of GraRED-HP3 through IR experiments.
 Finally, conclusions are presented in section 5.
@@
 Bredies et al.
-[5] introduced a linear, bounded, self-adjoint and positive semi-definite operator admissible preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
-The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates $(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$.
+[9] introduced the degenerate PPP framework with the positive semi-definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
+The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
 Definition 2.1.
@@
 $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$
-is single-valued, of full-domain, and Lipschitz continuous.
-Therefore, the preconditioned proximal point iteration is written into
+is single-valued, full-domain, and Lipschitz continuous.
+Therefore, the PPP iteration is written into
 (6) $\mathbf{u}^{0}\in\mathcal{H},\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}.$
-If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly non-expansive operator [1, section 23] and (6) becomes the standard proximal point iteration.
-If $\mathcal{M}$ is semi-definite, the property of $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [6, 9, 41, 5], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
+If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly nonexpansive (FNE)operator [2, section 23] and (6) becomes the standard proximal point iteration.
+If $\mathcal{M}$ is positive semi-definite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
 The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
@@
 $\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0,\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
-According to [5, Lemma 2.6], if $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
+The following lemma demonstrates that the firmly non-expansiveness notation can be generalized to the degenerate case [11, 15].
+If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
+Lemma 2.3 ([9]).
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone.
+Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
 (7) $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|_{\mathcal{M}}^{2}+\left\|(I-\mathcal{T})\mathbf{u}-(I-\mathcal{T})\mathbf{v}\right\|^{2}_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{v}\right\|^{2}_{\mathcal{M}}.$
-By the KM iteration, the relaxed PPP algorithm is written into
+By the KM iteration, the degenerate PPP method is written into
 (8) $\mathbf{u}^{k+1}=(1-\lambda_{k})\mathbf{u}^{k}+\lambda_{k}\mathcal{T}\mathbf{u}^{k},$
-where $\lambda_{k}$ satifies $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
-2.2 Implicit Gradient RED
-Romano et al. introduced an explicit regularizer called Regularization by Denoising (RED) [33] to construct an explicit objective function, which is defined by
-(9) $g(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
-where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser ($\sigma>0$ denotes a noise level).
-If $D_{\sigma}(\mathbf{x})$ satisfies homogeneity $D_{\sigma}(c\mathbf{x})=cD_{\sigma}(\mathbf{x})(\forall c>0)$ and has symmetric Jacobian, then $\nabla g(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})$.
-Under the following nonexpansive Assumption (A[35] on residul $R=I-D_{\sigma}$, i.e.,
-(A) $\left\|(\mathbf{x}-D_{\sigma}(\mathbf{x}))-(\mathbf{y}-D_{\sigma}(\mathbf{y}))\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n},$
-we will prove that there exists an implicit regularization $\phi$ such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
-Here is the definition of the proximal operator and characterization of proximal operators.
-Definition 2.3.
-[4, Section 6] Given a function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
+where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
+2.2 PnP and RED
+2.2.1 PnP prior
+Venkatakrishnan et al.
+proposed the first PnP method based on ADMM [56].
+PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$.
+PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60].
+Sreehari et al.
+established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51].
+When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$.
+Subsequently, a non-local means denoiser satisfying these conditions was constructed [51].
+Algorithm 1 PnP-ADMM
+1:Input: Given $\mathbf{u}^{0}=\mathbf{0},\mathbf{x}^{0},\mathbf{z}^{0}$ and total iterations $N>0$
+2:for $k=0,1,2,\ldots,N-1$ do
+3:     $\mathbf{x}^{k+1}=D_{\sigma}(\mathbf{z}^{k}-\mathbf{u}^{k})$
+4:     $\mathbf{z}^{k+1}=\text{prox}_{\lambda f}(\mathbf{x}^{k+1}+\mathbf{u}^{k})$
+5:     $\mathbf{u}^{k+1}=\mathbf{u}^{k}+(\mathbf{x}^{k+1}-\mathbf{z}^{k+1})$
+6:end for
+7:Output: $\mathbf{x}^{N}$
+Definition 2.4 ([5]).
+Given $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
 (9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{x}).$
@@
 A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c.
-(lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ if, and only if the following conditions hold jointly:
+(lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if, and only if the following conditions hold jointly:
   1. (a)
-    there exists a (convex l.s.c.) function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
+    there exists a convex l.s.c. function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
   2. (b)
@@
     $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
+2.2.2 RED prior
+Romano et al. introduced a famous RED [47], which is defined by
+(10) $g_{\text{red}}(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
+where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions,
+  1. (C1)
+    Local Homogeneity: For all $\mathbf{x}\in\mathbb{R}^{n}$, $D_{\sigma}\left((1+\epsilon)x\right)=(1+\epsilon)D_{\sigma}(\mathbf{x})$ for sufficiently small $\epsilon>0$.
+  2. (C2)
+    Differentiability: The denoiser $D_{\sigma}(\cdot)$ is differentiable.
+  3. (C3)
+    Jacobian Symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{n}$.
+  4. (C4)
+    Strong Passivity: The spectral radius of the Jacobian satisfies $\eta\left(\nabla D_{\sigma}(\mathbf{x})\right)\leq 1$.
+If $D_{\sigma}(\mathbf{x})$ satisfies C1C2, and C3, then $\nabla g_{\text{red}}(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})=R(\mathbf{x})$.
+Moreover, if the denoiser satisfies C4, then RED is convex and $D_{\sigma}$ is nonexpansive.
+2.2.3 Implicit Gradient RED
+The gradient of RED is exactly the denoising residual.
+We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function as proximal operators of implicit regularizers.
+Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
+The nonexpansive assumption has theoretically played a crucial role in PnP and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49],
+(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|,$
+for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual.
+Under the conditons C1C2C3 and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
 Lemma 2.6.
-Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, then the gradient of RED defines an implicit regularization $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ such that
-(11) $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}),D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}).$
-Proof 2.6.
-Let $\psi(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,h(\mathbf{x})=R(\mathbf{x})$ in Proposition 2.4.
-By Assumption (A), there exist a function denoted $\phi^{*}$($\phi^{*}$ is the conjugate of $\phi$) such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$.
-By Moreau decomposition [4, Theorem 6.44], $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$.
+Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditons C1C2C3C4 and assumption (A), then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
+(11) $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}).$
+Proof 2.7.
+By assumption condtions, the RED term $g_{\text{red}}$ is convex and $\nabla g_{\text{red}}=I-D_{\sigma}=R$.
+Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser.
+Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex.
+Applying Proposition 2.5 together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=I$, then there exist a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
+Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required.
+From the perspective of monotone operator theory, Lemma 2.8 [2, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
+Lemma 2.8.
+The mapping $T:\mathcal{X}\to\mathcal{X}$ is the resolvent of a maximal monotone operator if and only if $T$ is FNE.
+Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$.
+Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator.
+Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudo-contractive denoisers [57], have made the nonexpansive assumption more realistic in practice.
+Constructing the FNE denoiser $D_{\sigma}$ is feasible, and similarly, the residual $R=I-D_{\sigma}$ can also be FNE in (7) with $\mathcal{M}=I$.
+They can both be the proximal operators of implicit regularizers.
 3 Halpern-type preconditioned proximal point (HPPP)
-Compared to KM iteration, Halpern iteration [19] is an effective method to find a particular fixed point, i.e.,
-(12) $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k},$
-the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection of the initial point $\mathbf{u}^{0}$ to $\mathrm{Fix}(T)$  [19, 26, 40], i.e., $\mathbf{u}^{*}=P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$, where $P_{\Omega}(\mathbf{u}^{0})=\arg\min_{\mathbf{u}\in\Omega}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$ denotes the standard metric projection.
-Using the PPP to solve the inclusion problem (3), the fixed point of $\mathcal{T}$ is the solution of (3).
-Choosing an appropriate anchor point, such as the degraded image or an image filled with ones or zeros, can bring the projection closer to the ground-truth solution, resulting in better restoration performance.
-Hence, the HPPP algorithm (5) theoretically converges to the ground-truth solution under some known anchor point prior $\mathbf{a}$.
-When $\mathcal{T}$ satisfies the mild condition $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ [5], the HPPP algorithm is able to find a unique solution $\mathbf{u}^{*}$ of (3), where $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(see Lemma A.4).
+Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration [26]
+$\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$
+is an implicitly regularized method to find a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$.
+When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$.
+Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
 3.1 Convergence analysis
@@
 Theorem 3.1.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
 Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5).
@@
     $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0\text{ or }\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$.
-Then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$ which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
+Then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
 Proof 3.2.
-Firstly, we show that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$, where
-$\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}.$
-Then, we have
+First, we show $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$.
+Based on the definition of the upper limit, we take a subsequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ such that
+(12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
+By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$.
+According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
+Applying Lemma A.2 together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$, then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$.
+Hence, we have
+$\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},(\mathcal{T}\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{k_{n_{j}}})+(\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*})\right\rangle_{\mathcal{M}}$
+$=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
+$=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
+By Lemma B.1, we have
+(13) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0.$
+Denote $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, notice that
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}^{2}$
@@
 $\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+\mu_{k+1}\delta_{k+1},$
-where $\delta_{k}=\mu_{k}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2\mu_{k}(1-\mu_{k})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k-1}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$.
-By Lemma A.2, $\{\mathbf{u}^{k}\}$ is bounded, there exists $\tilde{\mathbf{u}}$ such that $\mathbf{u}^{k}\rightharpoonup\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$ according to the known condition.
-Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\tilde{\mathbf{u}}$ such that
-$\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
-$=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
-Since $\mathbf{u}^{*}$ is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{a}-\mathbf{u}\right\|_{\mathcal{M}}^{2}$ according to Lemma A.4, which solves
-$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
-Therefore $\limsup_{k\to\infty}\delta_{k}\leq 0$, from Lemma A.1 we obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$.
-we then prove $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$.
-$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
+applying Lemma A.1 together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, then we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
+By the definition of $\mathbf{u}^{k+1}$ we immediately obtain
+$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|$
 $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
-$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
-it easily follows that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$ from $\mu_{k+1}\to 0$ and $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$.
-It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfy conditions (i)-(iii).
-Under mild conditions(i)-(iii), compared with [2], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point.
-Corollary 3.3.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
-Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5).
-If $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=\infty,\mu_{k}\to 0(k\to\infty),\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$, which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
+(14) $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
+where the last inequality follows from (26) and $C$ is defined in (25).
+Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, and thus combined with (14) yields that
+$\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|=0,$
+which ends the proof of Theorem 3.1.
+It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfes conditions (i)-(iii) of Theorem 3.1.
+Compared with [3], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point.
+The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$.
+By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
+According to Theorem 3.1, we obtain the following Proposition 3.3.
+Proposition 3.3.
+If the conditions of Theorem 3.1 hold, and $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a maximal monotone operator, then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
 Proof 3.4.
-Assume that $\mathbf{u}^{k}\rightharpoonup\mathbf{u}$, $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ (Lemma A.2), it follows that
-$\mathcal{T}\mathbf{u}^{k}=\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}+\mathbf{u}^{k}\rightharpoonup 0+\mathbf{u}=\mathbf{u},$
-i.e., $\mathcal{T}\mathbf{u}^{k}\rightharpoonup\mathbf{u}$.
-From $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ and $\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}$ we have
-(13) $\mathcal{A}\mathcal{T}\mathbf{u}^{k}\owns\mathcal{M}(\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k})\to 0.$
-By the maximality of $\mathcal{A}$ we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ (see [1, Proposition 20.38]), hence $0\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}$ is a fixed point of $\mathcal{T}$.
-Thus, every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, it holds by Theorem 3.1.
+Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, it follows that
+$\mathcal{T}\mathbf{u}^{k_{n}}=(\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}})+\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{0}+\mathbf{u}=\mathbf{u},$
+i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$ we have
+$\mathcal{A}\mathcal{T}\mathbf{u}^{k_{n}}\owns\mathcal{M}(\mathbf{u}^{k_{n}}-\mathcal{T}\mathbf{u}^{k_{n}})\to\mathbf{0}.$
+Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$.
+Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
+This completes the proof of Proposition 3.3.
 Remark 3.5.
-Compared Theorem 3.1, Corollary 3.3 with [5, Theorem 2.9, convergence] and [5, Corollary 2.10], $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
-All conditions are the same except we only add the mild assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$ used for convergence analysis.
-$\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can converges weakly to the same fixed point of $\mathcal{T}$ since $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$.
+Compared Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
+All conditions are the same except for additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$.
 The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
+Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly, where $U$ is a closed convex cone and $V$ is a closed hyperplane [31].
+Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence.
+Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
+Counterexample 1.
+Suppose that $\mathcal{H}$ is infinite-dimensional and separable. Let $V$ and $U$ be Hundal’s hyperplane and Hundal’s cone, respectively, and set
+$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\in V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
+Then the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP (8) with $\lambda_{k}=1$ converges weakly, but not strongly, to a zero of $\mathcal{A}_{1}+\mathcal{A}_{2}$.
+Proof 3.6.
+As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the following sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
+(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k}$
+generated by the DRS algortihm can weakly converges to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
+Since the above (15) can be viewed as a special PPP with
+$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
+and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$.
+By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$.
+Indeed, if by contraction $\mathbf{u}^{k}\to\mathbf{0}$, then
+$\left\|\mathbf{w}^{k}-\mathbf{0}\right\|=\left\|\mathcal{C}^{*}\mathbf{u}^{k}-\mathbf{0}\right\|\leq\left\|\mathcal{C}\right\|\cdot\left\|\mathbf{u}^{k}-\mathbf{0}\right\|\to 0,$
+which contradicts the fact $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.
+Remark 3.7.
+From Counterexample 1, HPPP has a strong convergence advantage over PPP in infinite-dimensional spaces, which has important theoretical value.
+Remark 3.8.
+Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2).
+The restarted technique is an effective method in machine learning [41] and LP [17, 38].
+Algorithm 2 Restarted HPPP
+1:Initialization $\mathbf{u}^{0,0}=\mathbf{a}^{0}$, total iteration $N>0$, period $q>0$, and epoch $N_{e}=\lfloor\frac{N}{q}\rfloor$
+2:for $n=1,2,\cdots,N_{e}$ do
+3:     for $k=0,1,\cdots,q-1$ do
+4:         $\mathbf{u}^{n,k+1}=\mathrm{HPPP}(\mathbf{u}^{n,k},\mathbf{a}^{n})$
+5:     end for
+6:     $\mathbf{a}^{n+1}=\mathbf{u}^{n,q}$
+7:end for
+8:$\mathbf{u}^{N_{e},q}$
 3.2 Convergence rate
-In this section, we will give the result of the sub-linear convergence rate of the gap between two successive iterations. We now present a technical Lemma [36] for convergence rate.
+Based on the following technical Lemma, Sabach and Shtern [50] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
 Lemma 3.9 ([50]).
-Let $M>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers which satisfy $a_{1}<M$ and
+Let $M_{1}>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers such that $a_{1}<M_{1}$ and
 $a_{k+1}\leq(1-\gamma b_{k+1})a_{k}+(b_{k}-b_{k+1})c_{k},k\geq 1,$
-where $\gamma\in(0,1]$, $\{b_{k}\}_{k\in\mathbb{N}}$ is a sequence which is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M<\infty$.
-Then, the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
-$a_{k}\leq\frac{MJ}{\gamma k},k\geq 1,$
+where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$.
+Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
+$a_{k}\leq\frac{M_{1}J}{\gamma k},k\geq 1,$
 where $J=\lfloor\frac{2}{\gamma}\rfloor$.
-Corollary 3.7.
-Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{u^{k}\}$ be the sequence generated by (5).
+By applying the Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
+Proposition 3.10.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\{u^{k}\}$ be the sequence generated by (5).
 If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
   1. (i)
-    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2M^{\prime}}{k}$, i.e., $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$;
+    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$.
   2. (ii)
     $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
-Proof 3.8.
-For (i), denote $M^{\prime}=\sup_{k\in\mathbb{N}}\{\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}$, since
-$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M^{\prime},$
-it derives (i) by Lemma 3.6. For (ii), since
+Proof 3.11.
+(i)According to (30), we have
+$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
+where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined by (27).
+Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$ and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
+(ii)By (33), when $k\geq 2$ we have
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|$
-$\leq\frac{2M^{\prime}C}{k-1}+\frac{2M^{\prime}}{k(k+1)},k\geq 2.$
-and thus $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
-3.3 HPPP with TV regularization
-Chambolle-Pock(CP) is a first-order primal-dual algorithm to solve non-smooth convex optimization problems with known saddle-point structure [10].
-Take the starting point $(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$, the CP primal-dual algorithm is written into the following form:
-(14) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=(I+s\partial g^{*})^{-1}\left(\mathbf{y}_{k}+s\mathbf{K}(2\mathbf{x}_{k+1}-\mathbf{x}_{k})\right),\end{array}\right.$
-here we consider TV regularization $\mathbf{K}=\nabla$.
-The convergence of CP algorithm was proved when step sizes satisfy $\tau s\left\|\mathbf{K}\right\|^{2}<1$.
-By introducing the operator $\mathcal{A}$ and preconditioner $\mathcal{M}$, set
-(15) $\mathcal{A}=\begin{pmatrix}\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}.$
-As fistly noticed in [20, 5], the CP method is a special PPP method $0\in\mathcal{A}\mathbf{x}^{k+1}+\mathcal{M}(\mathbf{x}^{k+1}-\mathbf{x}^{k})$.
-Bredies et al.
-proved the weak convergence of the CP algorithm about the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ [5], i.e., $\mathcal{M}$ is a semi-definite preconditioner.
-To apply HPPP with TV regularization, under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$ we obtain
-(16) $\left\{\begin{array}[]{ll}\mathbf{x}_{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})(I+\tau\partial f)^{-1}(\mathbf{x}_{k}-\tau\mathbf{K}^{*}\mathbf{y}_{k}),\\ \mathbf{y}_{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\frac{2s\mathbf{K}(\mathbf{x}_{k+1}-\mu_{k+1}\mathbf{x}_{a})}{1-\mu_{k+1}}-s\mathbf{K}\mathbf{x}_{k}+\mathbf{y}_{k}\right),\end{array}\right.$
-where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points, respectively, and $\mu_{k}$ satisfies conditions(i)-(iii) of Theorem 3.1.
-3.4 HPPP with implicit regularization
-Recently, RED frameworks have achieved state-of-the-art performance for inverse problems by utilizing CNN (convolutional neural network) denoiser as regularization [14], such as DnCNN [46].
-Ryu et al.
-proposed a nonexpansive residual of DnCNN to ensure the convergence of PnP methods [35].
-Gradient step denoisers, defined as $D_{\sigma}=I-\nabla g_{\sigma}$ for Gaussian noise level $\sigma$, have been proposed to generalize RED [13, 24], where $g_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}$ is a scalar function parameterized by a differentiable neural network.
-The gradient step denoiser can be interpreted as the proximal operator of an implicit regularization [25].
-The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$ to solve (3).
-According to subsection 2.2, $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, we consider the special application $g=\phi,\mathbf{K}=I$ with the degenerate $\mathcal{M}(\tau s=1)$.
-The proposed HPPP algorithm has the advantage of converging to a particular solution than PPP.
-Due to the data adaptability of CNN, which has the powerful expressiveness to represent the implicit regularization $\phi$ [44, 45].
-Based on HPPP and GraRED, we propose the implicit Gradient RED via HPPP called GraRED-HP3 (see Algorithm 1).
-According to (8), we further discuss GraRED-P3 Algorithm 2 for IR problems.
-PnP-ADMM [39, 11, 35] is a well-known optimization method to solve (1) when $g=\phi$, i.e.,
-(17) $\mathbf{z}^{k+1}=$ $\mathrm{prox}_{\phi}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k})=D_{\sigma}(\mathbf{v}^{k}-\bar{\mathbf{u}}^{k}),$
-$\mathbf{v}^{k+1}=$ $\mathrm{prox}_{\lambda f}(\mathbf{z}^{k+1}+\bar{\mathbf{u}}^{k}),$
-$\bar{\mathbf{u}}^{k+1}=$ $\bar{\mathbf{u}}^{k}+\mathbf{z}^{k+1}-\mathbf{v}^{k+1}.$
-Let $\mathbf{w}^{k}=\mathbf{v}^{k}+\bar{\mathbf{u}}^{k}$, then PnP-ADMM can be written into equivalent Douglas-Rachford Splitting (DRS) form:
-(18) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k}).$
-Using $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\lambda_{k}=1$ in GraRED-P3 results exactly in the above DRS iteration.
-Thus, PnP-ADMM can be obtained from the perspective of PPP [6, 5], and its convergence follows by [5, Corollary 2.15].
-The following Theorem 3.9 verifies the convergence of Algorithm 2 and Algorithm 2.
-Theorem 3.9.
-Assume that a denoiser $D_{\sigma}:\mathcal{H}\to\mathcal{H}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, relaxing parameter $20<\inf_{k}\lambda_{k}\leq\sup_{k}\lambda_{k}<$ 2 and RED is convex, then $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 2 converges to some uncertain solution $\mathbf{x}^{*}$ of (1), while $\{\mathbf{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 converges to the particular solution $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of (3).
-Proof 3.10.
-Followed by [5, Corollary 2.10] and Corollary 3.3.
+$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}(\frac{1}{k}),$
+where the second inequality follows from (i) and the fact that $\mu_{k}-\mu_{k+1}=\frac{2}{k(k+1)}(k\geq 2)$, and $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, and $C,C_{1}$ are defined in (25) and (27), respectively.
+As for the second result, by the triangle inequality and $\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, for $k\geq 2$,
+$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|$ $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\left\|\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}\right\|$
+$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|$
+$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}(\frac{1}{k}).$
+Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3].
+By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE, let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, then HPPP is equivalent to the following reduced algorithm:
+(16) $\mathbf{w}^{0}=\mathcal{C}^{*}\mathbf{u}^{0},\mathbf{w}^{k+1}=\mu_{k+1}\mathbf{w}^{0}+(1-\mu_{k+1})\tilde{\mathcal{T}}\mathbf{w}^{k}.$
+By [36, Theorem 2.1] and [52, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
+Proposition 3.12.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16).
+If $\mu_{k}=\frac{1}{k+1}$ and $\mathbf{a}=\mathbf{u}^{0}$, then
+  1. (i)
+    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$.
+  2. (ii)
+    $\left\|\mathbf{u}^{k}-{\mathcal{T}}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}}{k+1}$ for $k\geq 0$ and $\mathbf{u}^{*}\in\mathrm{Fix}({\mathcal{T}})$.
+3.3 GraRED-HP3
+The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$.
+Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
+(17) $\left\{\begin{array}[]{ll}\mathbf{d}^{k}&=(I+\tau\lambda\partial f)^{-1}(\mathbf{x}^{k}-\tau\mathbf{K}^{*}\mathbf{y}^{k}),\\ \mathbf{x}^{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k},\\ \mathbf{y}^{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\mathbf{y}^{k}+s\mathbf{K}(2\mathbf{d}^{k}-\mathbf{x}^{k})\right),\end{array}\right.$
+where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points.
+Under the ideal conditions of RED and with Assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$.
+Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$, then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17).
+Set $\mathbf{K}=I$, the implicit Gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3.
+Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers.
+According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
+3.3.1 Acceleration for PnP methods
+PnP-ADMM (see Algorithm 1) is a well-known method to solve (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$.
+When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration.
+Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9].
+The related GraRED-P3 algorithm is shown in Algorithm 4.
+Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
+Remark 3.13.
+In [28], a worst-case convergence rate of $\mathcal{O}(\frac{1}{\sqrt{k}})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM.
+In contrast, GraRED-HP3 obtain an accelerated convergence rate of $\mathcal{O}(\frac{1}{k})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
 Algorithm 3 GraRED-HP3
-1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, diminishing anchor coefficient $\mu_{k}\in[0,1],\mu_{k}\to 0(k\to\infty),\sum_{k=0}^{\infty}\mu_{k}=+\infty$, and $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.
+1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, and $R=I-D_{\sigma}$ is the residual.
 2:for $k=0,1,2,\cdots,N-1$ do
 3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
-4:     $\mathbf{x}^{k+1}=\mu_{k}\mathbf{x}_{a}+(1-\mu_{k})\mathbf{d}^{k}$
-5:     $\mathbf{v}^{k+1}=R\left(s(2\mathbf{d}^{k}-\mathbf{x}^{k})+\mathbf{y}^{k}\right)$
-6:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})\mathbf{v}^{k+1}$
-7:end for
-8:$\mathbf{x}^{N}$.
+4:     $\mathbf{x}^{k+1}=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k}$
+5:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})R\left(\mathbf{y}^{k}+s(2\mathbf{d}^{k}-\mathbf{x}^{k})\right)$
+6:end for
+7:$\mathbf{x}^{N}$.
 Algorithm 4 GraRED-P3
-1:initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, $R=I-D_{\sigma}$ is a residual of denoiser $D_{\sigma}$, $\mathcal{A}=\begin{pmatrix}\lambda\partial f&{I}\\ -I&\partial{\phi}^{*}\end{pmatrix}$, degenerate preconditioner $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix},\tau s\leq 1$, where $I$ is identity matrix.
+1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, and $R=I-D_{\sigma}$ is the residual.
 2:for $k=0,1,2,\cdots,N-1$ do
@@
 7:$\mathbf{x}^{N}$.
+Algorithm 5 Relaxed RED-PRO via SD
+1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$ and the denoiser $T$.
+2:for $k=0,1,2,\ldots,N-1$ do
+3:     $\mathbf{x}^{k,0}=\mathbf{x}^{k}$
+4:     for $j=0,1,2,\ldots,J-1$ do
+5:         $\mathbf{x}^{k,j+1}=T_{\alpha}\left(t_{j}\mathbf{x}^{k}+(1-t_{j})\mathbf{x}^{k,j}\right)$
+6:     end for
+7:     $\mathbf{v}^{k}=\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\mathbf{x}^{k}+\left(1-\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\right)\mathbf{x}^{k,J}$
+8:     $\mathbf{x}^{k+1}=\mathbf{x}^{k}-\mu\left(\nabla f(\mathbf{x}^{k})+\lambda(\mathbf{x}^{k}-\mathbf{v}^{k})\right)$
+9:end for
+10:Output: $\mathbf{x}^{k+1}$
+3.3.2 Compared with RED-PRO
+Cohen et al. proposed the RED-PRO model [20],
+(18) $\min_{\mathbf{x}\in\mathrm{Fix}(T)}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}.$
+RED-PRO sheds new insights on denoiser prior, which theoretically bridges between RED and PnP prior.
+However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions.
+To address this issue, they relax the hard constraint of $\mathrm{Fix}(T)$ and replace it with a dilated fixed-point set, defined for some $\delta>0$ as
+$B_{\delta}(T)=\left\{\mathbf{x}\in\mathbb{R}^{n}:\left\|\mathbf{x}-P_{\mathrm{Fix}(T)}(\mathbf{x})\right\|\leq\delta\right\},$
+where $P_{\mathrm{Fix}(T)}(\mathbf{x})$ is the fixed-point projection of $\mathbf{x}$ onto $\mathrm{Fix}(T)$. The relaxed RED-PRO (RRP) problem is formulated as
+(19) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}+\frac{\lambda}{2}\left\|\mathbf{x}-P_{B_{\delta}(T)}(\mathbf{x})\right\|^{2}.$
+Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demi-contractive denoiser $T$.
+When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and the line 5 in RRP is equivalent to the Halpern iteration.
+Comparing Algorithm 3 with the inner loop of Algorithm 5, both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients.
+In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
+The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semi-definite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$.
+From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop.
+In contrast, $\mathcal{T}$ in GraRED-HP3 is an algorithmic operator and has no inner loop.
+From a PPP standpoint, we integrate HPPP with denoiser priors to propose Algorithm 3.
 4 Experiments
 In this section, we show the numerical experiments of the algorithms discussed in section 3.
-Firstly, we will verify the regularity of HPPP by an easy 1D example.
-Then, we compare the CP (14), PPP (8), GraRED-HP3 algorithms (see Algorithm 1), and GraRED-P3 (see Algorithm 2) for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms.
+Firstly, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and Restarted HPPP by other toy examples.
+Then, we compare the CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms.
+The source code is available at https://github.com/zsc15/HPPP.
 4.1 A toy example
@@
 (20) $\min_{x\in\mathbb{R}}f(x)+g(x),$
-where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$ (See Figure 1).
-Figure 1: Plot $f(x)+g(x)$.
-Since for any $y\in\mathbb{R}$,
-$\min\{(1+y)x-1,yx\}=\begin{cases}(1+y)x-1,&x<1\\ yx,&x\geq 1\end{cases},$
-then $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y),y\in\mathbb{R}$, where
-$\delta_{C}(y)=\begin{cases}0,&y\in C,\\ \infty,&y\notin C.\end{cases}$
-Therefore, the corresponding saddle-point problem is:
-$\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}xy+f(x)-g^{*}(y).$
+where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, we plot $f(x)+g(x)$ in Figure 1.
+[…]
+Figure 1: The image of $f(x)+g(x)$.
+The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$.
 We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$.
@@
 Fixed $x^{*}\geq 1$, then
-$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0;\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
+$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0,\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
 If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$ and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to contradiction.
 Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
-For this toy example shown in Figure 2, we choose the same initial point $\mathbf{u}^{0}=(-6,6)$ or $(0,0)$, three different anchor points $\mathbf{a}=(12,10),(12,9),(12,8)$, the semi-definite $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, and the total itertation number $N=1000$.
-As shown in Figure 2, $\mathbf{u}^{k}$ generated by HPPP (5) can converge to the particular saddle point controlled by the anchor point.
-While the limit of $\mathbf{u}^{k}$ generated by PPP (8) may be uncertain, which is related with $\mathbf{u}^{0}$ and $\lambda_{k}$.
-For example, the sequence $\mathbf{u}^{k}$ oscillates around the limit $(1.8,0)$ shown in Figure 2(d).
-Figure 2: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ in the $xOy$ Cartesian coordinate system.
-As shown in Figure 3, $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}=\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be geometrically interpreted as $\mathcal{M}$-projection, i.e., the point $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ projection onto the line $x-y=x_{a}-y_{a}$ through the anchor point $\mathbf{a}=(x_{a},y_{a})$.
-Therefore, the $\mathcal{M}$-projection of $\mathbf{a}=(x_{a},y_{a})$ onto $\Omega$ is definitely solved by
-$P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0;\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
+When $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, as shown in Figure 2, the minimization $\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}=\arg\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be interpreted as finding the $\mathcal{M}$-projection, namely, the projection of the anchor point $\mathbf{a}=(x_{a},y_{a})$ onto the line $x-y=x_{a}-y_{a}$.
+Therefore, the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\Omega$ is uniquely determined,
+(21) $P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0,\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
+$x-y-1\leq 0$$x$$y$$O$$S(2,0)$$R(3,1)$
 Figure 2: $\mathcal{M}$-projection onto the saddle point set.
+We set the total iteration number $N=1000$.
+As shown in Figure 3(a), with the anchor point fixed at $\mathbf{a}=(12,9)$, the sequence $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP converges to the same point $\mathbf{u}^{*}=(3,0)$, regardless of the initial point, which verifies Proposition 3.3 and (21).
+HPPP is also an implicitly regularized method.
+In contrast, Figure 3(b) shows that the sequence generated by PPP oscillates around the limit $(1.8,0)$.
+However, a quantitative mathematical characterization of the limit of the PPP sequence remains elusive.
+Figure 3: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP or PPP.
+4.2 A 2D toy example
+We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41], for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
+$T_{\theta}\mathbf{x}=\frac{1}{\gamma}\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\mathbf{x}.$
+We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$) and Restarted HPPP ($q=50$).
+Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$.
+Figure 4 shows that HPPP and Restarted HPPP indeed provide accelerated residual decay.
+Figure 4: Trajectories, distance to solution, two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
+4.3 An infinite-dimensional example
+We consider an infinite dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences.
+For any $\mathbf{x}=(x_{0},x_{1},x_{2},\cdots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$.
+We consider the following operator inclusion problem:
+(22) $\mathbf{0}\in\mathcal{A}_{1}\mathbf{x}+\mathcal{A}_{2}\mathbf{x},\mathcal{A}_{1}=-\mathcal{S}-\mathbf{b},\mathcal{A}_{2}=I,$
+where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\cdots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\cdots)$.
+Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, then $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29].
+Additionally, by solving the above inclusion problem (22) we obtain $\mathbf{x}^{*}=(1,0,\cdots)$.
+Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semi-definite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$.
+We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\cdots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$.
+For this example, we present visual results about the residual and the distance to the solution.
+As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
+(e) Distance to solution
+(f) Residual comparison
+Figure 5: Comparison with PPP to solve an infinite dimensional problem in $\ell^{2}$ space.
 4.4 Image deblurring
-Firstly, we compared with classic TV regularization about CP [10], PPP [5] , and HPPP(see details in 3.3). For classic TV -$\ell^{2}$ regularization,
+Firstly, we compared with CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
 (23) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|_{2}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
-where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator.
-In case $\mathbf{A}\mathbf{x}$ can be written as a convolution, i.e., $\mathbf{A}\mathbf{x}=\mathbf{k}\ast\mathbf{x}$, where $\mathbf{k}$ is the blurring convolution kernel.
-The primal-dual model of TV -$\ell^{2}$ deblurring is
-(21) $\min_{\mathbf{u}\in\mathbb{R}^{n}}\max_{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{u},\mathrm{div}\mathbf{p}\rangle+\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2}-\delta_{P}(\mathbf{p}),$
-where $f(\mathbf{u})=\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{u}-\mathbf{y}\right\|_{2}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm defined as:
-$\left\|\mathbf{p}\right\|_{\infty}=\max_{i,j}|\mathbf{p}_{i,j}|,|\mathbf{p}_{i,j}|=\sqrt{(\mathbf{p}_{i,j}^{1})^{2}+(\mathbf{p}_{i,j}^{2})^{2}}.$
-We can calculate the proximal operator $(I+\tau\partial g^{*})^{-1}$ of the indicator $g^{*}(\mathbf{p})$, i.e.,
-$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}.$
-The resolvent operator for $f(\mathbf{x})$ can be computed by FFT.
-$\mathbf{x}$ $=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})$
-$=\arg\min_{\mathbf{x}}\frac{\|\mathbf{x}-\tilde{\mathbf{x}}\|}{2\tau}+\frac{\lambda}{2}\left\|\mathbf{k}*\mathbf{x}-\mathbf{y}\right\|_{2}^{2}$
-$=\mathcal{F}^{-1}\left(\frac{\tau\lambda\mathcal{F}(\mathbf{y})\mathcal{F}^{*}\left(\mathbf{k}\right)+\mathcal{F}(\tilde{\mathbf{u}})}{\tau\lambda\mathcal{F}\left(\mathbf{k}\right)^{2}+1}\right),$
-where $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the FFT and inverse FFT, respectively.
-We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive WGN with noise level $0.01$.
+where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator, and $\lambda,\beta$ are balance coefficients.
+We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$.
 Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP.
 All the algorithms use the degraded images as initial points.
-We calculate the norm $\left\|K\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$.
+We calculate the norm $K=\left\|\nabla\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$.
 Their parameters are given in Table 1.
@@
   • GraRED-HP3 : $\tau=1,s=1,\mu_{k}=1/(k+2),\lambda=20,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
-Secondly, we compared with RED and RED-PRO.
-The RED-PRO model uses the hybrid steepest descent method (HSD) [42, 43] to solve IR problems.
-Following [33, 14], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
-We finally obtained the degraded images with an additive WGN with noise level $\sigma=\sqrt{2}$.
-The original RGB image is converted to the YCbCr image, PnP restoration algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image.
-PSNR is measured on the luminance channel of the ground truth and the restored images.
 Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3.
-The performance of three different methods is evaluated using PSNR measure.
+The performance of these methods is evaluated using PSNR measure.
 The best recovery results are highlighted in bold.
-From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularization is more powerful to regularize inverse imaging problems.
+From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularizer is more powerful to regularize inverse imaging problems.
 We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6.
 To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization.
-As shown in Figure 5, we plot their respective evolutions of PSNR values for iterations for the image House with 10 random initial points.
-The HPPP algorithm converges faster (less than 200 iterations) and achieves better PSNR values than CP and PPP algorithms.
-Once the anchor point is chosen, the proposed algorithm is more robust than CP and PPP algorithms for image deblurring with random initializations.
-To verify Corollary 3.7, we show the trend of convergence rate of the gap $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ between two successive iterations throughout the iterations with step size $\min\{\frac{2}{k},1\}(k\geq 1)$ in Figure 6.
+As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point.
+Once the anchor point is chosen, the proposed HPPP algorithm is robust to random perturbations of the initial points, which achieves stable recovery.
 Table 1: Parameters on CP, PPP, and HPP with TV regularization for image deblurring.
@@
 PPP 26.05 30.85 25.99 27.53 27.85 25.51 26.88 24.46 28.98
-HPPP(TV) 26.00 31.39 26.05 27.65 27.99 25.51 26.82 24.51 29.09
+HPPP 26.00 31.39 26.05 27.65 27.99 25.51 26.82 24.51 29.09
 GraRED-P3 26.87 32.51 28.34 29.16 29.70 26.90 27.95 24.66 29.68
@@
 Refer to captionRefer to caption
-(f) HPPP ($27.65$dB)
+(f) HPPP ($27.99$dB)
 Refer to captionRefer to caption
@@
 (h) GraRED-HP3 ($29.67$dB)
-Figure 4: Deblurring result of Butterfly degraded with the Gaussian blur kernel with noise level $\sigma=0.01$.
-Figure 5: Random initialization.
-Figure 6: The convergence rate of $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ of GraRED-H3. The image Starfish is degraded by a Gaussian PSF with noise level $0.01$.
+Figure 6: Deblurring results of Butterfly degraded by the Gaussian blur kernel with noise level $\sigma=2.55$.
+(a) Butterfly
+(b) Parrots
+Figure 7: Robustness comparison with random initialization. Test images are degraded by the Gaussian blurring kernel with the noise level $\sigma=2.55$.
+Secondly, we compared with RED and RED-PRO.
+Following [47, 20], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
+We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$.
+The original RGB image is converted to the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image.
+PSNR is measured on the luminance channel of the ground truth and the restored images.
+Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3.
 From the deblurring experiment Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective.
-Table 3: Recovery results are obtained by RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3 with TNRD [12].
+We further compare the differences among GraRED-P3, GraRED-HP3, and Restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$.
+In Table 5, we report the average PSNR values of 15 test images using six different blur kernels.
+When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks.
+Moreover, by dynamically updating the anchor points, Restarted HPPP further boosts the performance of GraRED-HP3.
+(a) Recovery peroformance
+(b) Fixed point residual
+Figure 8: Recovery peroformance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
+Additionally, we trained an FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO.
+We also compared with state-of-the-art methods using advanced denoisers such as DRUNet ($D_{2}$[65] and GS denoiser ($D_{3}$[32, 33, 59].
+For GraRED-HP3, the anchor point is the same as the initial point.
+The default parameters are presented in Table 4.
+As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods.
+In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$).
+The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
+Table 3: Recovery results are obtained by different algorithms with TNRD [18]. The best two results are highlighted in red and blue colors, respectively.
 Algorithms Uniform kernel Gaussian kernel($\sigma_{k}=1.6$)
@@
 GraRED-HP3 26.80 30.88 30.74 32.42 28.06 31.80 32.27 33.50
-Figure 7: Deblurring result of Butterfly degraded by uniform kernel.
+Table 4: Parameters on GraRED-HP3 and Restarted HPPP with advanced denoisers.
+[…]
+Restarted HPPP$(D_{2})$ $\tau=1.2,s=1/1.2,N=30,q=10$
+[…]
+[…]
+Table 5: Compared with GraRED-P3, GraRED-HP3, and Restarted HPPP with TNRD for different tasks.
+kernels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+noise level, scale $\sigma=5,\times 3$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$
+GraRED-P3 27.49 29.25 30.92 28.19 33.53 33.73
+GraRED-HP3 27.61 29.33 30.99 28.32 33.53 33.71
+Restarted HPPP 27.66 29.36 31.00 28.34 33.53 33.73
+Table 6: Compared with state-of-the-art methods.
+Test gray images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$.
+The best two results are highlighted in red and blue colors, respectively.
+Pepper Craft Cameraman Couple Man House Starfish Butterfly Boat
+RED [47] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
+RED-PRO [20] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
+DPIR [65] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
+PnP-DRS [33] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
+PnP-PGD [32] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
+PnP-DYS [59] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
+[…]
+Restarted HPPP$(D_{2})$ 28.34 26.75 27.02 29.71 30.44 32.69 28.91 29.28 30.01
+[…]
+Restarted HPPP$(D_{3})$ 29.81 27.13 27.49 30.04 30.80 33.15 30.09 30.53 30.37
+Figure 9: Deblurring results of Butterfly degraded by the uniform kernel.
 4.5 Image inpainting
-In this section, we use the proposed algorithm to solve TV image inpainting problems and compare their numerical results with CP [10], PPP [5], and HPPP algorithms.
+In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [16], PPP [9], and HPPP algorithms.
 The discrete image inpainting model is
 $\min_{\mathbf{x}\in\mathbb{R}^{n}}\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
-where $\left\|\nabla\mathbf{x}\right\|_{1}$ is the TV regularization, $\lambda,\beta$ are balance coefficients, $\odot$ indicates pointwise multiplication. Then the saddle-point problem is
-$\min_{\mathbf{x}\in\mathbb{R}^{n}}\max_{p\in\mathbb{R}^{n}\times\mathbb{R}^{n}}-\langle\mathbf{x},\mathrm{div}\mathbf{p}\rangle+\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}-\delta_{P}(\mathbf{p}),$
-where $f(\mathbf{x})=\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2},g^{*}(\mathbf{p})=\delta_{P}(\mathbf{p}),P=\{\mathbf{p}\in\mathbb{R}^{n}\times\mathbb{R}^{n}:\left\|\mathbf{p}\right\|_{\infty}\leq\beta\}$, $\left\|\mathbf{p}\right\|_{\infty}$ is the discrete maximum norm.
-Their resolvent operators of $f,g^{*}$ are
-$\mathbf{p}=\left(I+\tau\partial g^{*}\right)^{-1}(\tilde{\mathbf{p}})\quad\Longleftrightarrow\quad\mathbf{p}_{i,j}=\frac{\tilde{\mathbf{p}}_{i,j}}{\max\left(\beta,\left|\tilde{\mathbf{p}}_{i,j}\right|\right)}$
-and
-$\mathbf{x}=(I+\tau\partial f)^{-1}(\tilde{\mathbf{x}})=\frac{2\tau\mathbf{M}\odot\mathbf{y}+\tilde{\mathbf{x}}}{1+2\tau\mathbf{M}},$
-the multiplication and division operators should be understood pointwise in the above formula.
+where $\mathbf{y}$ is the degraded image, $\mathbf{M}$ is a mask, and $\lambda,\beta$ are balance coefficients.
 We test 10 common images for evaluation.
@@
 The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed.
-All the algorithms start their iterations with the degraded images.
+All the algorithms start with the degraded image.
 For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters:
-  • HPPP: $\tau=s=1/\left\|K\right\|=0.57$, anchor point $\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$, stepsize $\mu_{k}=\frac{1}{10(k+2)}$;
+  • HPPP: $\tau=s=1/\left\|K\right\|=0.57,\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a},\mu_{k}=\frac{1}{10(k+2)}$;
   • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$ or $\lambda_{k}=1.2$;
@@
 In Table 7 and Table 8, we compared the numerical performance of classic algorithms with TV regularization.
-As we can see from both two tables, the proposed two algorithms outperform other algorithms.
-In Figure 8 and 9, we compare visualization results of House degraded by Bernoulli random mask and character mask, the proposed algorithms achieve better visual performance than TV regularization.
-Moreover, we compare the recovery results about different anchors $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})$ with fixed $\mathbf{y}_{a}=\mathbf{0}$ and step size $\frac{1}{k+1}$.
-As shown in Figure 10, the anchor $\mathbf{x}_{a}=\mathbf{1}$ can achieve the best peroformance for random inpainting, which illustates the projection point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ of $\mathbf{a}=(\mathbf{1},\mathbf{0})$ is closest to the true solution.
-Anchors selection is not difficult, GraRED-HP3 can achieve similar peroformance with other anchors.
-Table 4: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, Bernoulli random missing).
+In Figure 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask, the proposed algorithms achieve better visual performance than TV regularization.
+Table 7: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, Bernoulli random missing).
 Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
@@
 PPP 23.58 29.05 24.53 24.32 23.76 23.43 23.33 23.13 26.36
-HPPP(TV) 23.89 29.19 24.55 24.46 24.04 23.52 23.44 23.35 26.45
+HPPP 23.89 29.19 24.55 24.46 24.04 23.52 23.44 23.35 26.45
 GraRED-P3 29.92 36.15 32.67 31.99 31.95 29.53 30.05 32.25 32.91
 GraRED-HP3 29.90 36.14 32.68 31.97 31.92 29.52 30.04 32.23 32.90
-Table 5: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=0.01$, character texture missing).
+Table 8: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, character texture missing).
 Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
@@
 (h) GraRED-HP3 ($36.14$dB)
-Figure 8: Recovery results of House degraded by random mask.
+Figure 10: Recovery results of House degraded by the random mask.
 Refer to captionRefer to caption
-(a) Clean image
+(i) Clean image
 Refer to captionRefer to caption
-(b) Degraded image
+(j) Degraded image
 Refer to captionRefer to caption
-(c) CP ($31.58$dB)
+(k) CP ($31.58$dB)
 Refer to captionRefer to caption
-(d) $\lambda_{k}=1.6$: PPP ($31.79$dB)
+(l) $\lambda_{k}=1.6$: PPP ($31.79$dB)
 Refer to captionRefer to caption
-(e) $\lambda_{k}=1.2$: PPP ($31.74$dB)
+(m) $\lambda_{k}=1.2$: PPP ($31.74$dB)
 Refer to captionRefer to caption
-(f) HPPP ($32.00$dB)
+(n) HPPP ($32.00$dB)
 Refer to captionRefer to caption
-(g) GraRED-P3 ($37.46$dB)
+(o) GraRED-P3 ($37.46$dB)
 Refer to captionRefer to caption
-(h) GraRED-HP3 ($37.47$dB)
-Figure 9: Recovery results of House degraded by character mask.
-(a) Clean image
-(b) Degraded image
-(c) Anchor 1: 31.89dB
-(d) Anchor 2: 31.95dB
-(e) Anchor 3: 31.97dB
-(f) Anchor 4: 31.85dB
-Figure 10: Compared with different anchors for random inpainting, i.e., $\mathbf{x}_{a}=\mathbf{0},\mathbf{M}\odot\mathbf{x}_{0}+0.5(1-\mathbf{M}),\mathbf{1},\mathbf{x}_{0}$.
+(p) GraRED-HP3 ($37.47$dB)
+Figure 11: Recovery results of House degraded by the character mask.
 5 Conclusions
-In this paper, based on Halpern iteration and gradient RED, we propose HPPP and GraRED-HP3 for IR problems, which can converge strongly to a particular fixed point $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
-Numerical experiments verify the regularity of HPPP, and the effectiveness GraRED-HP3 for image deblurring and inpainting, which can achieve better performance than classic algorithms with TV regularization.
-In the future, we plan to study the convergence of nonconvex implicit regularization [25] and extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [3].
-Appendix A Boundedness and $\mathcal{M}$-projection
+This paper introduces a Halpern-type PPP (HPPP) algorithm, which has the advantage of strong convergence over PPP.
+In addition, GraRED-HP3 is proposed by integrating HPPP with denoiser priors for IR problems, which is an accelerated PnP-ADMM.
+Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments.
+In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [4].
+Acknowledgments
+The authors thank Prof.
+Defeng Sun for bringing the interesting work [52] to our attention and Mr.
+Feng Xue for his insightful discussions.
+We sincerely appreciate Associate Editor Michael Elad and the anonymous referees for their thorough review and valuable comments, which have substantially improved this paper.
+Appendix A Boundedness
 Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5).
-To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important Lemma from [23, Lemma 2.5].
+To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important lemma from [30, Lemma 2.5].
 Lemma A.1.
@@
 Lemma A.2.
-Let $\mathcal{T}$ be a $\mathcal{M}$-FNE operator, $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz, and $\{\mathbf{u}^{k}\}$ be the sequence generated by (5), $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$, and $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, which satisfies the following:
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5).
+Then the following assertions are hold:
   1. (i)
-    $\{\mathbf{u}^{k}\},\{\mathcal{T}\mathbf{u}^{k}\}(k\in\mathbb{N})$ are bounded;
+    The sequences $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ are bounded;
   2. (ii)
-    $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0,k\to\infty$;
+    If the sequence $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$;
   3. (iii)
-    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0,\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0,k\to\infty$.
+    If the conditions of (ii) hold, then $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|=0$.
 Proof A.3.
-Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$, $C=L\left\|\mathcal{C}\right\|$. For $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$,
+Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [9, Proposition 2.3]. Denote
+(25) $C=L\left\|\mathcal{C}\right\|,$
+for $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$, we have
 $\left\|\mathcal{T}\mathbf{u}^{\prime}-\mathcal{T}\mathbf{u}^{\prime\prime}\right\|$ $=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime}-(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime\prime}\right\|$
-$\leq L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
-For (i), $\mathbf{u}^{k}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$,
+$=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}(\mathcal{C}^{*}\mathbf{u}^{\prime}-\mathcal{C}^{*}\mathbf{u}^{\prime\prime})\right\|$
+$\leq L\left\|\mathcal{C}\right\|\left\|\mathcal{C}^{*}(\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime})\right\|$
+$=L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
+(26) $=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}},$
+where the third inequality follows from the $L$-Lipschitz property of $(\mathcal{M}+\mathcal{A})^{-1}$, and the fourth equality follows from the definition of $\left\|\mathbf{u}\right\|_{\mathcal{M}}$, i.e., $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}=\left\|\mathcal{C}^{*}\mathbf{u}\right\|$ for any $\mathbf{u}\in\mathcal{H}$.
+First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$, we have
 $\left\lVert\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\rVert_{\mathcal{M}}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}$
@@
 $\leq\dots\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}.$
+Denote
+(27) $C_{1}=\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\},$
+then we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq C_{1}$ for all $k\in\mathbb{N}$.
 Furthermore,
@@
 $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\cdot C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
-$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}<+\infty.$
-and
-$\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}<+\infty$
-So is the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
-For (ii), we should show $\mathcal{T}$ is $\mathcal{M}$-asymptotically regular $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$.
-$\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$. There exists a positive real number $M>0$ such that
-$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq M,$
-it follows from the $\mathcal{M}$-FNE operator $\mathcal{T}$ and boundedness of $\{\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\}_{k\in\mathbb{N}}$,
-(23) $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
+$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,CC_{1}\}<+\infty.$
+By (26), we have
+(28) $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq CC_{1}<+\infty.$
+So the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is also bounded.
+For (ii), we will prove $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$.
+Let $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}$, by the triangle inequality and the $\mathcal{M}$-non-expansiveness of $\mathcal{T}$, we have
+$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$
+$\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
+(29) $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}=M.$
+Note that $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
+$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
 $\leq(1-\mu_{k+1})\left\|\mathcal{T}u^{k}-\mathcal{T}u^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$
-$\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|M$
-If $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$ from Lemma A.1.
-(24) $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}+\mathbf{u}^{k+1}-\mathcal{T}u^{k}\right\|_{\mathcal{M}}$
-$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
-Thus, $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}\to 0$.
-For (iii), there exists $M^{\prime}$ such that $\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|\leq M^{\prime}$,
-$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{u}^{0}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
-$\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
-Therefore $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|\to 0$.
-$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\to 0,k\to\infty.$
-Lemma A.4.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$, and $\mathrm{Fix}(\mathcal{T})$ be a closed convex subset of a Hilbert space $\mathcal{H}$, and $\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function.
-There exists a unique solution $u^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
-(25) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
-Proof A.5.
-Let $l(\mathbf{u})=\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is the differentiable convex function.
-Assume that $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ is the optimal solution, $\mathrm{Fix}(\mathcal{T})$ is the convex set, thus $t\in(0,1),\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for $\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$,
+(30) $\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
+where the last inequality follows from Lemma 2.3 and (29).
+Applying Lemma A.1 together with $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$ to (30), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$.
+Since $\mathbf{u}^{k+1}-\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, we have
+$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mathbf{u}^{k}-\mathbf{u}^{k+1})+(\mathbf{u}^{k+1}-\mathcal{T}{\mathbf{u}^{k}})\right\|_{\mathcal{M}}$
+(31) $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
+Using (29) and $\lim_{k\to\infty}\mu_{k}=0$ in (31), we can obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$.
+For (iii), let $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, by the triangle inequality and (28), we have
+(32) $\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}=M^{\prime}.$
+Similar to the proof of (30), we have
+$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
+(33) $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
+where the last inequality follows from (26).
+Applying (30) and $\lim_{k\to\infty}\mu_{k}=0$ to the above inequality (33), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0$.
+Similar to the proof of (31),
+$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|.$
+By $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\mu_{k}=0$ and boundedness of $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$, we finally prove $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=0$.
+Appendix B $\mathcal{M}$-projection
+Lemma B.1.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
+$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
+Proof B.2.
+Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function.
+Assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$.
+Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$.
+Hence,
 $\lim_{t\to 0}\frac{l(\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*}))-l(\mathbf{u}^{*})}{t}$ $=\left\langle l^{\prime}(\mathbf{u}^{*}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
-$=\left\langle 2\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
-$=2\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
-If $\mathbf{u}^{**}$ is the another solution such that $\left\langle\mathbf{u}^{**}-a,\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0$.
-Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in the above two inequalities, respectively.
-(26) $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
-(27) $\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
-If we add two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. Since
-$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$
-It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$ from above inequality.
-As mentioned in [2, 9] and uniqueness projection onto $\mathrm{Fix}(\mathcal{T})$, we can introduce the following notion of $\mathcal{M}$-projection.
-Definition A.6 ($\mathcal{M}$-projection).
-Assume $\forall\mathbf{u}^{0}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{u}^{0}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{u}^{0}\right\|_{\mathcal{M}}(\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}))$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}-$projection of $\mathbf{u}^{0}$ onto $\mathrm{Fix}(\mathcal{T})$, also denoted $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$.
-Lemma A.7.
-Let $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$ such that $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ and $\mathbf{u}^{0}\in\mathcal{H}$, then the following conditions are equivalent:
+$=\left\langle\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
+(34) $=\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
+If $\mathbf{u}^{**}$ is the another solution, then it also satisfies
+(35) $\left\langle\mathbf{u}^{**}-{\mathbf{a}},\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
+Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in (34) and (35), respectively, then
+$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
+$\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
+Add above two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
+$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0.$
+It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$, which completes the proof of Lemma B.1.
+As mentioned in [3, 15], we can introduce the following notion of $\mathcal{M}$-projection.
+Definition B.3 ($\mathcal{M}$-projection).
+Assume $\mathbf{a}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
+The following Lemma B.4 extends the variational inequality of the metric projection.
+Lemma B.4.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Then the following assertions are equivalent:
   1. (i)
-    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{u}^{0})$;
+    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$;
   2. (ii)
-    $\left\langle\mathbf{u}^{*}-\mathbf{u}^{0},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
-Proof A.8.
-See Lemma A.4 and Definition A.6.
+    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
+Proof B.5.
+See Lemma B.1 and Definition B.3.

### v5

@@
 Abstract
-Recently, the degenerate preconditioned proximal point (PPP) method provides a unified and flexible framework for designing and analyzing operator-splitting algorithms such as Douglas-Rachford (DR).
+Recently, the degenerate preconditioned proximal point (PPP) method provided a unified and flexible framework for designing and analyzing operator-splitting algorithms, such as the Douglas-Rachford (DR) splitting.
 However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants.
 To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method.
-Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors such as Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method.
-Finally, numerical experiments including several toy examples and image restoration validate the effectiveness of our proposed algorithms.
+Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors, such as the Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method.
+Finally, numerical experiments, including several toy examples and image restoration, validate the effectiveness of our proposed algorithms.
 keywords:
@@
 Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces.
-The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization such as TV [48] to mitigate the ill-posedness of IR problems.
+The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization, such as TV (total variation) [48], to mitigate the ill-posedness of IR problems.
 By the first-order optimality condition, the convex optimization problem (1) is equivalent to the following inclusion problem:
@@
 where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3].
-Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as:
+Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as
 […]
@@
 The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain.
-Another way to solve (1) is to transform it into the saddle-point problem [16, 43], i.e.,
+Another way to solve (1) is to transform it into a saddle-point problem [16, 43], i.e.,
 (4) $\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}\left\langle\mathbf{K}\mathbf{x},\mathbf{y}\right\rangle+\lambda f(\mathbf{x})-g^{*}(\mathbf{y}),$
-where $g^{*}:\mathcal{K}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$.
-In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) splitting method [16, 43] and the Primal-Dual Hybrid Gradient (PDHG) method [27], are designed and studied extensively.
+where $g^{*}:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$.
+In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) primal-dual method [16, 43] and the primal-dual hybrid gradient (PDHG) method [27], are designed and studied extensively.
 In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
 Recently, Bredies et al.
-[11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semi-definite $\mathcal{M}$.
-By choosing appropriate preconditioners, this framework could cover the DR and CP splitting algorithms [9].
+[11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semidefinite $\mathcal{M}$.
+By choosing appropriate preconditioners, this framework could cover the DR and CP algorithms [9].
 It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34].
@@
 Similarly, Chen et al.
-introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semi-proximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38].
+introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semiproximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38].
 A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8].
@@
 Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction.
-Based on the Halpern iteration, the Halpern-type preconditioned proximal point algorithm called HPPP (5) is thus proposed to overcome the two limitations of the degenerate PPP method,
+Based on the Halpern iteration, the following Halpern-type preconditioned proximal point algorithm (called HPPP) (5) is thus proposed to overcome the two limitations of the degenerate PPP method:
 (5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k},$
-where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
+where $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$, and $\mathcal{A},\mathcal{M}$ are detailed in subsection 3.3, and $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
 PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53].
@@
 provided a consensus equilibrium interpretation on denoiser priors [14].
-Romano et al.
-introduced RED (Regularization by Denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization.
-Powerful denoisers (e.g., DnCNN [66]) typically do not meet the conditions of RED [45].
+Romano , Elad, and Milanfar introduced RED (regularization by denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization.
+Powerful denoisers such as denoising convolutional neural network (DnCNN) [66] typically do not meet the conditions of RED [45].
 To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45].
@@
 Meanwhile, Ryu et al.
-proved the convergence of PnP-FBS and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49].
+proved the convergence of PnP-FBS (forward-backward splitting) and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49].
 However, enforcing strong convexity on $f$ precludes many IR tasks [33].
 Furthermore, Cohen et al.
-parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault et al.
-proposed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33].
-Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on their equivalence [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
-Based on HPPP and denoiser priors,  [^1] Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.we propose the Gradient Regularization by Denoising via HPPP called GraRED-HP3 for IR problems.
-The main contributions are as follows:
+parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault, Leclaire, and Papadakis developed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33].
+Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on the equivalence between ADMM and DRS [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
+Based on HPPP and PnP priors, we propose the gradient regularization by denoising via HPPP called GraRED-HP3 (Algorithm 3111Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual..
+Our main contributions are as follows:
   1. 1.
     Theoretical contributions.
-The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semi-definite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution.
+The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semidefinite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution.
 Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
@@
 We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint.
-Noting that the PnP-ADMM algorithm can be reformulated as an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
+Noting that the PnP-ADMM algorithm can be reformulated in an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
   3. 3.
     Experimental validation.
-We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency with several toy examples.
+We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency, with several toy examples.
 In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
@@
 Bredies et al.
-[9] introduced the degenerate PPP framework with the positive semi-definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
+[9] introduced the degenerate PPP framework with the positive semidefinite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
 The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
 Definition 2.1.
-An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bouned, linear, self-adjoint, and positive semi-definite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
+An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bounded, linear, self-adjoint, and positive semidefinite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
 $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$
 is single-valued, full-domain, and Lipschitz continuous.
-Therefore, the PPP iteration is written into
+Therefore, the PPP iteration is written as
 (6) $\mathbf{u}^{0}\in\mathcal{H},\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}.$
-If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly nonexpansive (FNE)operator [2, section 23] and (6) becomes the standard proximal point iteration.
-If $\mathcal{M}$ is positive semi-definite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
+If $\mathcal{M}=I$, then $\mathcal{T}$ is a firmly nonexpansive (FNE) operator [2, Proposition 23.8] and (6) becomes the standard proximal point iteration.
+If $\mathcal{M}$ is positive semidefinite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
 The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
 Definition 2.2 ($\mathcal{M}$-monotonicity).
-Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bouned linear positive semi-definite operator, then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
-$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0,\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
-The following lemma demonstrates that the firmly non-expansiveness notation can be generalized to the degenerate case [11, 15].
+Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bounded linear positive semidefinite operator; then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
+$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0\quad\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
+The following lemma demonstrates that the firmly nonexpansive notation can be generalized to the degenerate case [11, 15].
 If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
 Lemma 2.3 ([9]).
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone.
 Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
 (7) $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|_{\mathcal{M}}^{2}+\left\|(I-\mathcal{T})\mathbf{u}-(I-\mathcal{T})\mathbf{v}\right\|^{2}_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{v}\right\|^{2}_{\mathcal{M}}.$
-By the KM iteration, the degenerate PPP method is written into
+By the KM iteration, the degenerate PPP method is written as
 (8) $\mathbf{u}^{k+1}=(1-\lambda_{k})\mathbf{u}^{k}+\lambda_{k}\mathcal{T}\mathbf{u}^{k},$
-where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
-2.2 PnP and RED
+where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k\in\mathbb{N}}\lambda_{k}(2-\lambda_{k})=+\infty$.
+2.2 PnP prior and RED
 2.2.1 PnP prior
-Venkatakrishnan et al.
-proposed the first PnP method based on ADMM [56].
+Venkatakrishnan, Bouman, and Wohlberg proposed the first PnP method based on ADMM [56].
 PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$.
@@
 When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$.
-Subsequently, a non-local means denoiser satisfying these conditions was constructed [51].
+Subsequently, a nonlocal means denoiser satisfying these conditions was constructed [51].
 Algorithm 1 PnP-ADMM
@@
 Definition 2.4 ([5]).
-Given $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
-(9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{x}).$
+The proximal operator of $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined by
+(9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{u}\in\mathbb{R}^{n}}\{\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{u})\}.$
 Proposition 2.5 ([40, 25]).
-A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c.
-(lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if, and only if the following conditions hold jointly:
+A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex lower semicontinuous function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if and only if the following conditions hold jointly:
   1. (a)
-    there exists a convex l.s.c. function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
+    There exists a convex lower semicontinuous function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
   2. (b)
     $h$ is nonexpansive, i.e.,
-    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
+    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|\quad\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
 2.2.2 RED prior
-Romano et al. introduced a famous RED [47], which is defined by
+RED [47] is defined by
 (10) $g_{\text{red}}(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
-where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions,
+where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions:
   1. (C1)
@@
   3. (C3)
-    Jacobian Symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{n}$.
+    Jacobian symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
   4. (C4)
@@
 The gradient of RED is exactly the denoising residual.
-We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function as proximal operators of implicit regularizers.
+We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function are proximal operators of implicit regularizers.
 Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
-The nonexpansive assumption has theoretically played a crucial role in PnP and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49],
-(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|,$
+The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49]:
+(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|;$
 for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual.
-Under the conditons C1C2C3 and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
+Under the conditions C1C2C3, and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and we prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
 Lemma 2.6.
-Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditons C1C2C3C4 and assumption (A), then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
+Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditions C1C2C3C4 and assumption (A); then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
 (11) $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}).$
 Proof 2.7.
-By assumption condtions, the RED term $g_{\text{red}}$ is convex and $\nabla g_{\text{red}}=I-D_{\sigma}=R$.
+By assumption conditions, the RED term $g_{\text{red}}$ is convex, and $\nabla g_{\text{red}}=I-D_{\sigma}=R$.
 Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser.
 Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex.
-Applying Proposition 2.5 together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=I$, then there exist a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
+Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
 Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required.
@@
 Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator.
-Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudo-contractive denoisers [57], have made the nonexpansive assumption more realistic in practice.
-Constructing the FNE denoiser $D_{\sigma}$ is feasible, and similarly, the residual $R=I-D_{\sigma}$ can also be FNE in (7) with $\mathcal{M}=I$.
-They can both be the proximal operators of implicit regularizers.
+Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudocontractive denoisers [57], have made the nonexpansive assumption more realistic in practice.
+Constructing the FNE denoiser $D_{\sigma}$ is feasible, similarly, the residual $R=I-D_{\sigma}$ can be FNE in (7) with $\mathcal{M}=I$.
+Both can be proximal operators of implicit regularizers.
 3 Halpern-type preconditioned proximal point (HPPP)
-Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration [26]
-$\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$
-is an implicitly regularized method to find a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$.
+Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29].
+The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [26] is an implicitly regularized method for finding a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$.
 When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$.
@@
 Theorem 3.1.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5).
-Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
+Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and that $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
   1. (i)
-    $\lim_{k\to\infty}\mu_{k}=0$;
+    $\lim_{k\to\infty}\mu_{k}=0$,
   2. (ii)
-    $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty$;
+    $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty$,
   3. (iii)
@@
 (12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
-By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$.
+By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, and we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$.
 According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
-Applying Lemma A.2 together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$, then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$.
+Applying Lemma A.2, together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$; then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$.
 Hence, we have
@@
 (13) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0.$
-Denote $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, notice that
+Denoting $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, noticing that
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}^{2}$
@@
 $\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+\mu_{k+1}\delta_{k+1},$
-applying Lemma A.1 together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, then we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
+and applying Lemma A.1, together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
 By the definition of $\mathbf{u}^{k+1}$ we immediately obtain
@@
 (14) $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
-where the last inequality follows from (26) and $C$ is defined in (25).
-Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, and thus combined with (14) yields that
+where the last inequality follows from (26) and $C$ is defined as in (25).
+Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, combined with (14) they yield that
 $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|=0,$
 which ends the proof of Theorem 3.1.
-It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfes conditions (i)-(iii) of Theorem 3.1.
-Compared with [3], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point.
+It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfies conditions (i)-(iii) of Theorem 3.1.
+Compared with [3], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point.
 The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$.
 By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
-According to Theorem 3.1, we obtain the following Proposition 3.3.
+According to Theorem 3.1, we obtain the following proposition.
 Proposition 3.3.
@@
 Proof 3.4.
-Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, it follows that
+Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, and it follows that
 $\mathcal{T}\mathbf{u}^{k_{n}}=(\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}})+\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{0}+\mathbf{u}=\mathbf{u},$
-i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$ we have
+i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, we have
 $\mathcal{A}\mathcal{T}\mathbf{u}^{k_{n}}\owns\mathcal{M}(\mathbf{u}^{k_{n}}-\mathcal{T}\mathbf{u}^{k_{n}})\to\mathbf{0}.$
-Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$.
+Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$.
 Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
@@
 Remark 3.5.
-Compared Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
-All conditions are the same except for additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$.
-The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
-Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly, where $U$ is a closed convex cone and $V$ is a closed hyperplane [31].
+We compare Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
+All conditions are the same, except for the additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$.
+The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption, especially in applications to splitting algorithms, and is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
+Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [31], where $U$ is a closed convex cone and $V$ is a closed hyperplane.
 Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence.
@@
 Suppose that $\mathcal{H}$ is infinite-dimensional and separable. Let $V$ and $U$ be Hundal’s hyperplane and Hundal’s cone, respectively, and set
-$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\in V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
+$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\notin V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
 Then the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP (8) with $\lambda_{k}=1$ converges weakly, but not strongly, to a zero of $\mathcal{A}_{1}+\mathcal{A}_{2}$.
 Proof 3.6.
-As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the following sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
-(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k}$
-generated by the DRS algortihm can weakly converges to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
-Since the above (15) can be viewed as a special PPP with
-$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
+As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
+(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k},$
+generated by the DRS algorithm can weakly converge to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
+The sequence (15) can be viewed as a special PPP with
+$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\quad\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
 and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$.
 By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$.
-Indeed, if by contraction $\mathbf{u}^{k}\to\mathbf{0}$, then
+Indeed, if by contraction, $\mathbf{u}^{k}\to\mathbf{0}$, then
 $\left\|\mathbf{w}^{k}-\mathbf{0}\right\|=\left\|\mathcal{C}^{*}\mathbf{u}^{k}-\mathbf{0}\right\|\leq\left\|\mathcal{C}\right\|\cdot\left\|\mathbf{u}^{k}-\mathbf{0}\right\|\to 0,$
-which contradicts the fact $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.
+which contradicts the fact that $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.
 Remark 3.7.
@@
 1:Initialization $\mathbf{u}^{0,0}=\mathbf{a}^{0}$, total iteration $N>0$, period $q>0$, and epoch $N_{e}=\lfloor\frac{N}{q}\rfloor$
-2:for $n=1,2,\cdots,N_{e}$ do
-3:     for $k=0,1,\cdots,q-1$ do
+2:for $n=1,2,\ldots,N_{e}$ do
+3:     for $k=0,1,\ldots,q-1$ do
 4:         $\mathbf{u}^{n,k+1}=\mathrm{HPPP}(\mathbf{u}^{n,k},\mathbf{a}^{n})$
@@
 $a_{k+1}\leq(1-\gamma b_{k+1})a_{k}+(b_{k}-b_{k+1})c_{k},k\geq 1,$
-where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$.
+where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined as in $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$.
 Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
 where $J=\lfloor\frac{2}{\gamma}\rfloor$.
-By applying the Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
+By applying Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed-point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
 Proposition 3.10.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Let $\{u^{k}\}$ be the sequence generated by (5).
@@
   1. (i)
-    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$.
+    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}\left(\frac{1}{k}\right)$;
   2. (ii)
-    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
+    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}\left(\frac{1}{k}\right)$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}\left(\frac{1}{k}\right)$.
 Proof 3.11.
@@
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
-where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined by (27).
-Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$ and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
+where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined as in (27).
+Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$, and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
 (ii) By (33), when $k\geq 2$ we have
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|$
-$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}(\frac{1}{k}),$
+$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}\left(\frac{1}{k}\right),$
 where the second inequality follows from (i) and the fact that $\mu_{k}-\mu_{k+1}=\frac{2}{k(k+1)}(k\geq 2)$, and $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, and $C,C_{1}$ are defined in (25) and (27), respectively.
@@
 $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|$
-$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}(\frac{1}{k}).$
+$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}\left(\frac{1}{k}\right).$
 Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3].
-By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE, let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, then HPPP is equivalent to the following reduced algorithm:
+By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE.
+Let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, and then HPPP is equivalent to the following reduced algorithm:
 (16) $\mathbf{w}^{0}=\mathcal{C}^{*}\mathbf{u}^{0},\mathbf{w}^{k+1}=\mu_{k+1}\mathbf{w}^{0}+(1-\mu_{k+1})\tilde{\mathcal{T}}\mathbf{w}^{k}.$
@@
 Proposition 3.12.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16).
@@
   1. (i)
-    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$.
+    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$,
   2. (ii)
@@
 3.3 GraRED-HP3
-The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$.
+The primal-dual algorithm for solving (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$.
 Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
@@
 Under the ideal conditions of RED and with assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$.
-Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$, then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17).
-Set $\mathbf{K}=I$, the implicit Gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3.
+Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17).
+Set $\mathbf{K}=I$; the implicit gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3.
 Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers.
@@
 3.3.1 Acceleration for PnP methods
-PnP-ADMM (see Algorithm 1) is a well-known method to solve (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$.
+PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$.
 When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration.
@@
 Remark 3.13.
-In [28], a worst-case convergence rate of $\mathcal{O}(\frac{1}{\sqrt{k}})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM.
-In contrast, GraRED-HP3 obtain an accelerated convergence rate of $\mathcal{O}(\frac{1}{k})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
+In [28], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM.
+In contrast, GraRED-HP3 obtains an accelerated convergence rate of $\mathcal{O}\left(\frac{1}{k}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
 Algorithm 3 GraRED-HP3
 1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, and $R=I-D_{\sigma}$ is the residual.
-2:for $k=0,1,2,\cdots,N-1$ do
+2:for $k=0,1,2,\ldots,N-1$ do
 3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
@@
 1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, and $R=I-D_{\sigma}$ is the residual.
-2:for $k=0,1,2,\cdots,N-1$ do
+2:for $k=0,1,2,\ldots,N-1$ do
 3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
@@
 Algorithm 5 Relaxed RED-PRO via SD
-1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$ and the denoiser $T$.
+1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$, and the denoiser $T$.
 2:for $k=0,1,2,\ldots,N-1$ do
@@
 9:end for
-10:Output: $\mathbf{x}^{k+1}$
+10:Output: $\mathbf{x}^{N}$
 3.3.2 Compared with RED-PRO
-Cohen et al. proposed the RED-PRO model [20],
+RED-PRO [20] is defined by
 (18) $\min_{\mathbf{x}\in\mathrm{Fix}(T)}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}.$
-RED-PRO sheds new insights on denoiser prior, which theoretically bridges between RED and PnP prior.
+RED-PRO sheds new insights on the denoiser prior, which theoretically bridges the gap between RED and the PnP prior.
 However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions.
@@
 (19) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}+\frac{\lambda}{2}\left\|\mathbf{x}-P_{B_{\delta}(T)}(\mathbf{x})\right\|^{2}.$
-Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demi-contractive denoiser $T$.
-When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and the line 5 in RRP is equivalent to the Halpern iteration.
-Comparing Algorithm 3 with the inner loop of Algorithm 5, both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients.
+Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demicontractive denoiser $T$.
+When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and line 5 in RRP is equivalent to the Halpern iteration.
+Comparing Algorithm 3 with the inner loop of Algorithm 5, we see that both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients.
 In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
-The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semi-definite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$.
+The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semidefinite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$.
 From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop.
@@
 In this section, we show the numerical experiments of the algorithms discussed in section 3.
-Firstly, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and Restarted HPPP by other toy examples.
-Then, we compare the CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms.
+First, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and restarted HPPP by other toy examples.
+Then, we compare CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting and verify the efficiency of the proposed algorithms.
 The source code is available at https://github.com/zsc15/HPPP.
 4.1 A toy example
-We consider the following optimization problem in $\mathbb{R}$, i.e.,
+We consider the optimization problem in $\mathbb{R}$, i.e.,
 (20) $\min_{x\in\mathbb{R}}f(x)+g(x),$
-where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, we plot $f(x)+g(x)$ in Figure 1.
+where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, and we plot $f(x)+g(x)$ in Figure 1.
 […]
@@
 We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$.
-Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$.
-Fixed $x^{*}\geq 1$, then
+Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*})\text{ for all }(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$.
+Fix $x^{*}\geq 1$; then
 $\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0,\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
-If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$ and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to contradiction.
+If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction.
 Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@
 $x-y-1\leq 0$$x$$y$$O$$S(2,0)$$R(3,1)$
-Figure 2: $\mathcal{M}$-projection onto the saddle point set.
+Figure 2: $\mathcal{M}$-projection onto the saddle-point set.
 We set the total iteration number $N=1000$.
@@
 4.2 A 2D toy example
-We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41], for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
+We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
 $T_{\theta}\mathbf{x}=\frac{1}{\gamma}\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\mathbf{x}.$
-We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$) and Restarted HPPP ($q=50$).
+We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$), and restarted HPPP ($q=50$).
 Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$.
 Figure 4 shows that HPPP and restarted HPPP indeed provide accelerated residual decay.
-Figure 4: Trajectories, distance to solution, two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
+Figure 4: Trajectories, distance to solution, and two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
 4.3 An infinite-dimensional example
-We consider an infinite dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences.
-For any $\mathbf{x}=(x_{0},x_{1},x_{2},\cdots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$.
+We consider an infinite-dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences.
+For any $\mathbf{x}=(x_{0},x_{1},x_{2},\ldots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$.
 We consider the following operator inclusion problem:
 (22) $\mathbf{0}\in\mathcal{A}_{1}\mathbf{x}+\mathcal{A}_{2}\mathbf{x},\mathcal{A}_{1}=-\mathcal{S}-\mathbf{b},\mathcal{A}_{2}=I,$
-where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\cdots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\cdots)$.
-Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, then $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29].
-Additionally, by solving the above inclusion problem (22) we obtain $\mathbf{x}^{*}=(1,0,\cdots)$.
-Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semi-definite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$.
-We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\cdots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$.
+where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\ldots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\ldots)$.
+Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29].
+Additionally, by solving the above inclusion problem (22), we obtain $\mathbf{x}^{*}=(1,0,\ldots)$.
+Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semidefinite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$.
+We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\ldots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$.
 For this example, we present visual results about the residual and the distance to the solution.
@@
 (f) Residual comparison
-Figure 5: Comparison with PPP to solve an infinite dimensional problem in $\ell^{2}$ space.
+Figure 5: Comparison with PPP to solve an infinite-dimensional problem in $\ell^{2}$ space.
 4.4 Image deblurring
-Firstly, we compared with CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
+First, we compared CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
 (23) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|_{2}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
 where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator, and $\lambda,\beta$ are balance coefficients.
-We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$.
+We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test grayscale images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$.
 Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP.
@@
 Their parameters are given in Table 1.
-Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
+Both GraRED-P3 and GraRED-HP3 use DnCNN; other parameters are used as follows:
   • GraRED-P3 : $\tau=1,s=1,\lambda_{k}=0.2,\lambda=20$;
   • GraRED-HP3 : $\tau=1,s=1,\mu_{k}=1/(k+2),\lambda=20,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
-Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3.
+Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3, and GraRED-HP3.
 The performance of these methods is evaluated using PSNR measure.
 The best recovery results are highlighted in bold.
-From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularizer is more powerful to regularize inverse imaging problems.
+From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates that the implicit regularizer is more powerful for regularizing inverse imaging problems.
 We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6.
-To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization.
+We further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization.
 As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point.
@@
 HPPP $\tau=s=1/K=0.57,\mu_{k}=\frac{1}{k+2},\mathbf{x}_{a}=\mathbf{A}^{\mathrm{T}}\mathbf{y},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$
-Table 2: Deblurring results of gray images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
+Table 2: Deblurring results of grayscale images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
 Cameraman House Pepper Starfish Butterfly Craft Parrots Barbara Boat
@@
 Figure 7: Robustness comparison with random initialization. Test images are degraded by the Gaussian blurring kernel with the noise level $\sigma=2.55$.
-Secondly, we compared with RED and RED-PRO.
-Following [47, 20], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
+Second, we compared RED and RED-PRO.
+Following [47, 20], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
 We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$.
-The original RGB image is converted to the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image.
+The original RGB image is converted into the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image.
 PSNR is measured on the luminance channel of the ground truth and the restored images.
 Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3.
-From the deblurring experiment Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective.
+From the deblurring experiment of Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective.
 We further compare the differences among GraRED-P3, GraRED-HP3, and restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$.
@@
 Moreover, by dynamically updating the anchor points, restarted HPPP further boosts the performance of GraRED-HP3.
-(a) Recovery peroformance
+(a) Recovery performance
 (b) Fixed point residual
-Figure 8: Recovery peroformance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
-Additionally, we trained an FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO.
-We also compared with state-of-the-art methods using advanced denoisers such as DRUNet ($D_{2}$[65] and GS denoiser ($D_{3}$[32, 33, 59].
+Figure 8: Recovery performance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
+Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO.
+We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[65] and GS ($D_{3}$[32, 33].
+The PnP-DYS method [59] uses the Davis-Yin splitting (DYS) algorithm.
 For GraRED-HP3, the anchor point is the same as the initial point.
@@
 The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
-Table 3: Recovery results are obtained by different algorithms with TNRD [18]. The best two results are highlighted in red and blue colors, respectively.
+Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [18]. The best two results are highlighted in red and blue, respectively.
 Algorithms Uniform kernel Gaussian kernel($\sigma_{k}=1.6$)
@@
 […]
-Table 5: Compared with GraRED-P3, GraRED-HP3, and Restarted HPPP with TNRD for different tasks.
+Table 5: Comparison with GraRED-P3, GraRED-HP3, and restarted HPPP with TNRD for different tasks.
 Kernels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
 Restarted HPPP 27.66 29.36 31.00 28.34 33.53 33.73
-Table 6: Compared with state-of-the-art methods.
-Test gray images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$.
-The best two results are highlighted in red and blue colors, respectively.
+Table 6: Comparison with state-of-the-art methods.
+Test grayscale images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$.
+The best two results are highlighted in red and blue, respectively.
 Pepper Craft Cameraman Couple Man House Starfish Butterfly Boat
@@
 We test 10 common images for evaluation.
-The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed.
+The first $\mathbf{M}$ is filled with a Bernoulli random mask, where each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed.
 The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed.
 All the algorithms start with the degraded image.
-For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters:
+For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$ and use the following additional parameters:
   • HPPP: $\tau=s=1/\left\|K\right\|=0.57,\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a},\mu_{k}=\frac{1}{10(k+2)}$;
-  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$ or $\lambda_{k}=1.2$;
+  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$, or $\lambda_{k}=1.2$;
   • CP: $\tau=s=1/\left\|K\right\|=0.57$.
-Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used below:
+Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used:
   • GraRED-P3 : $\tau=10,s=0.1,\lambda_{k}=0.2,\lambda=5$;
   • GraRED-HP3 : $\tau=10,s=0.1,\mu_{k}=0.05/(k+2),\lambda=5,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
-In Table 7 and Table 8, we compared the numerical performance of classic algorithms with TV regularization.
-In Figure 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask, the proposed algorithms achieve better visual performance than TV regularization.
+In Tables 7 and 8, we compared the numerical performance of classic algorithms with TV regularization.
+In Figures 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask; the proposed algorithms achieve better visual performance than TV regularization.
 Table 7: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, Bernoulli random missing).
@@
 Appendix A Boundedness
-Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5).
-To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important lemma from [30, Lemma 2.5].
+First, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5).
+To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [30, Lemma 2.5].
 Lemma A.1.
@@
   1. (i)
-    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivently $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
+    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivalently, $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
   2. (ii)
@@
 Lemma A.2.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5).
@@
 (29) $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}=M.$
-Note that $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
+Note that by $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
 $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
@@
 Lemma B.1.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
-$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
+$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0\quad\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
 Proof B.2.
-Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function.
-Assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$.
+Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function, assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$.
 Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$.
@@
 $\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
-Add above two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
+Add the above two inequalities, we then obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
 $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0.$
@@
 Definition B.3 ($\mathcal{M}$-projection).
-Assume $\mathbf{a}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
+Assume $\mathbf{a}\in\mathcal{H}$ and that there exists a unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
 The following Lemma B.4 extends the variational inequality of the metric projection.
 Lemma B.4.
-Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz.
+Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz.
 Then the following assertions are equivalent:
@@
   2. (ii)
-    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
+    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
 Proof B.5.

### v6

@@
 1 Introduction
-Image restoration (IR) problems, including image deblurring, super-resolution, and inpainting, can be formulated as the following optimization problem [16, 27]:
+Image restoration (IR) problems, including image deblurring, super-resolution, and inpainting, can be formulated as the following optimization problem [Chambolle2011, He2012]:
 (1) $\min_{\mathbf{x}\in\mathcal{X}}\lambda f(\mathbf{x})+g(\mathbf{K}\mathbf{x}),$
@@
 Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces.
-The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization, such as TV (total variation) [48], to mitigate the ill-posedness of IR problems.
+The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization, such as TV (total variation) [RUDIN1992259], to mitigate the ill-posedness of IR problems.
 By the first-order optimality condition, the convex optimization problem (1) is equivalent to the following inclusion problem:
-[…]
-where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3].
-Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as
-[…]
+(2) $\text{find }\mathbf{x}\in\mathcal{X}\text{ such that }\mathbf{0}\in\lambda\partial f(\mathbf{x})+\mathbf{K}^{*}\partial g(\mathbf{K}\mathbf{x}),$
+where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [beck2017first, Chapter 3].
+Following [He2012, Bredies2022], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as
+(3) $\text{find }\mathbf{u}\in\mathcal{H}\text{ such that }\mathbf{0}\in\mathcal{A}\mathbf{u},$
 where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$.
-The problem (3) is common in modern optimization and variational analysis [2, 3].
-When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39].
-The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46].
+The problem (3) is common in modern optimization and variational analysis [BCombettes, Bauschke2023].
+When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [Minty1962MonotoneO].
+The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [Rockafellar1976].
 However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue.
-The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain.
-Another way to solve (1) is to transform it into a saddle-point problem [16, 43], i.e.,
+The well-known Douglas-Rachford splitting (DRS) [Eckstein1992] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain.
+Another way to solve (1) is to transform it into a saddle-point problem [Chambolle2011, Pock2011], i.e.,
 (4) $\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}\left\langle\mathbf{K}\mathbf{x},\mathbf{y}\right\rangle+\lambda f(\mathbf{x})-g^{*}(\mathbf{y}),$
 where $g^{*}:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$.
-In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) primal-dual method [16, 43] and the primal-dual hybrid gradient (PDHG) method [27], are designed and studied extensively.
-In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
+In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) primal-dual method [Chambolle2011, Pock2011] and the primal-dual hybrid gradient (PDHG) method [He2012], are designed and studied extensively.
+In particular, He and Yuan [He2012] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
 Recently, Bredies et al.
-[11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semidefinite $\mathcal{M}$.
-By choosing appropriate preconditioners, this framework could cover the DR and CP algorithms [9].
-It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34].
+[Bredies2015, Bredies2017, Bredies2022] developed a unified degenerate PPP algorithmic framework with a positive semidefinite $\mathcal{M}$.
+By choosing appropriate preconditioners, this framework could cover the DR and CP algorithms [Bredies2022].
+It is crucial to study algorithms that exhibit both strong convergence and acceleration [2002Iterative, BOT2021, Izuchukwu2023, Bauschke2023, Sun2024, chen2024hprlpimplementationhprmethod, lu2024restartedhalpernpdhglinear], such as strongly convergent proximal point methods [2002Iterative] and forward-reflected-backward splitting algorithms [Izuchukwu2023].
 For example, Bauschke et al.
-demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3].
+demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [Bauschke2023].
 In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point.
 Moreover, Sun et al.
-[52] proposed an accelerated preconditioned alternating direction method of multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6].
+[Sun2024] proposed an accelerated preconditioned alternating direction method of multipliers (pADMM) by leveraging the degenerate PPP method [Bredies2022] and the fast Krasnosel’skiĭ-Mann (KM) iteration [fastKM2023].
 Similarly, Chen et al.
-introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semiproximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38].
-A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8].
+introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semiproximal terms for efficiently solving linear programming (LP) problems [chen2024hprlpimplementationhprmethod], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [lu2024restartedhalpernpdhglinear].
+A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [BOT2021].
 Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces.
 Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
-The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29].
-Then the Halpern iteration is also known as an implicitly regularized method [22].
+The classic Halpern iteration [halpern1967fixed] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [he2024convergence].
+Then the Halpern iteration is also known as an implicitly regularized method [Diakonikolas2020].
 Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems.
-Achieving stable reconstruction for ill-posed inverse problems is important [24].
-Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29].
-Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction.
+Achieving stable reconstruction for ill-posed inverse problems is important [engl1996regularization].
+Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [TranDinh2024], which is widely utilized in machine learning [Diakonikolas2020, Park2022, he2024convergence].
+Notably, the work [Park2022] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction.
 Based on the Halpern iteration, the following Halpern-type preconditioned proximal point algorithm (called HPPP) (5) is thus proposed to overcome the two limitations of the degenerate PPP method:
@@
 where $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$, and $\mathcal{A},\mathcal{M}$ are detailed in subsection 3.3, and $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
-PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53].
+PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [Venkatakrishnan2013, PnP2015, PnP2020forMRI, PnP-CI2023, There-Operator-Splitting2024] and have achieved state-of-the-art performance in inverse imaging tasks [zhang2021plug, hurault2022proximal, Tan2024].
 Buzzard et al.
-provided a consensus equilibrium interpretation on denoiser priors [14].
-Romano , Elad, and Milanfar introduced RED (regularization by denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization.
-Powerful denoisers such as denoising convolutional neural network (DnCNN) [66] typically do not meet the conditions of RED [45].
-To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45].
+provided a consensus equilibrium interpretation on denoiser priors [Buzzard2018].
+Romano , Elad, and Milanfar introduced RED (regularization by denoising) [Romano2017], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [cohen2021regularization] that can be exploited in first-order optimization.
+Powerful denoisers such as denoising convolutional neural network (DnCNN) [Zhang2017] typically do not meet the conditions of RED [REDClarifications2019].
+To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [REDClarifications2019].
 Based on the fixed-point projection, Cohen et al.
-proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model.
+proposed the RED-PRO model [cohen2021regularization] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [yamada2001hybrid, Yamada2005] was employed to solve the resulting model.
 Meanwhile, Ryu et al.
-proved the convergence of PnP-FBS (forward-backward splitting) and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49].
-However, enforcing strong convexity on $f$ precludes many IR tasks [33].
+proved the convergence of PnP-FBS (forward-backward splitting) and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [ryu2019plug].
+However, enforcing strong convexity on $f$ precludes many IR tasks [hurault2022proximal].
 Furthermore, Cohen et al.
-parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault, Leclaire, and Papadakis developed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33].
-Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on the equivalence between ADMM and DRS [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
+parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [NEURIPS2021_97108695], and Hurault, Leclaire, and Papadakis developed the gradient step (GS) denoiser [hurault2022gradient, hurault2022proximal], which can be interpreted as a proximal operator of an implicit regularizer [hurault2022proximal].
+Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on the equivalence between ADMM and DRS [ryu2019plug, hurault2022proximal], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
 Based on HPPP and PnP priors, we propose the gradient regularization by denoising via HPPP called GraRED-HP3 (Algorithm 3111Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual..
@@
 The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semidefinite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution.
-Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
+Moreover, compared with the asymptotic regularity result of PPP [Bredies2022, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
   2. 2.
@@
 We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint.
-Noting that the PnP-ADMM algorithm can be reformulated in an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
+Noting that the PnP-ADMM algorithm can be reformulated in an equivalent DRS form [ryu2019plug, hurault2022proximal], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [He2015], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
   3. 3.
@@
 2 Preliminaries
-[…]
+In this section, we provide some fundamental concepts related to the degenerate PPP method and denoiser priors.
+Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and corresponding induced norm $\left\|\cdot\right\|$, and let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a (maybe multivalued) operator (that through the rest of this paper we often identify with its graph in $\mathcal{H}\times\mathcal{H}$).
 2.1 Preconditioned proximal point
 Bredies et al.
-[9] introduced the degenerate PPP framework with the positive semidefinite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
+[Bredies2022] introduced the degenerate PPP framework with the positive semidefinite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$.
 The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
@@
 $\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$
-is single-valued, full-domain, and Lipschitz continuous.
+is single-valued and has full domain.
 Therefore, the PPP iteration is written as
 (6) $\mathbf{u}^{0}\in\mathcal{H},\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}.$
-If $\mathcal{M}=I$, then $\mathcal{T}$ is a firmly nonexpansive (FNE) operator [2, Proposition 23.8] and (6) becomes the standard proximal point iteration.
-If $\mathcal{M}$ is positive semidefinite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
+If $\mathcal{M}=I$, then $\mathcal{T}$ is a firmly nonexpansive (FNE) operator [BCombettes, Proposition 23.8] and (6) becomes the standard proximal point iteration.
+If $\mathcal{M}$ is positive semidefinite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [Bredies2015, BUI2020124315, Xue2022, Bredies2022], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$.
 The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
@@
 $\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0\quad\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
-The following lemma demonstrates that the firmly nonexpansive notation can be generalized to the degenerate case [11, 15].
+The following lemma demonstrates that the firmly nonexpansive notation can be generalized to the degenerate case [Bredies2015, BUI2020124315].
 If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
-Lemma 2.3 ([9]).
+Lemma 2.3 ([Bredies2022]).
 Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone.
@@
 2.2.1 PnP prior
-Venkatakrishnan, Bouman, and Wohlberg proposed the first PnP method based on ADMM [56].
+Venkatakrishnan, Bouman, and Wohlberg proposed the first PnP method based on ADMM [Venkatakrishnan2013].
 PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$.
-PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60].
+PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [PnP-CI2023, wu2024principled].
 Sreehari et al.
-established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51].
+established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [PnP2015].
 When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$.
-Subsequently, a nonlocal means denoiser satisfying these conditions was constructed [51].
+Subsequently, a nonlocal means denoiser satisfying these conditions was constructed [PnP2015].
 Algorithm 1 PnP-ADMM
@@
 7:Output: $\mathbf{x}^{N}$
-Definition 2.4 ([5]).
+Definition 2.4 ([beck2017first]).
 The proximal operator of $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined by
 (9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{u}\in\mathbb{R}^{n}}\{\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{u})\}.$
-Proposition 2.5 ([40, 25]).
+Proposition 2.5 ([moreau1965proximite, Gribonval2020]).
 A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex lower semicontinuous function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if and only if the following conditions hold jointly:
@@
 2.2.2 RED prior
-RED [47] is defined by
+RED [Romano2017] is defined by
 (10) $g_{\text{red}}(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
@@
   3. (C3)
-    Jacobian symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
+    Jacobian symmetry [REDClarifications2019]: $\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
   4. (C4)
@@
 Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
-The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49]:
+The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [ryu2019plug]:
 (A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|;$
@@
 Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex.
-Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
+Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [beck2017first, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
 Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required.
-From the perspective of monotone operator theory, Lemma 2.8 [2, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
+From the perspective of monotone operator theory, Lemma 2.8 [BCombettes, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
 Lemma 2.8.
 The mapping $T:\mathcal{X}\to\mathcal{X}$ is the resolvent of a maximal monotone operator if and only if $T$ is FNE.
-Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$.
+Moreover, by [BCombettes, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$.
 Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator.
-Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudocontractive denoisers [57], have made the nonexpansive assumption more realistic in practice.
+Recent developments, such as spectral normalization [ryu2019plug, pmlr-v202-delattre23a], learning maximally monotone operators [MMO2021], learning FNE denoisers [building_FNE, bredies2024learningfirmlynonexpansiveoperators], and learning pseudocontractive denoisers [Wei2024], have made the nonexpansive assumption more realistic in practice.
 Constructing the FNE denoiser $D_{\sigma}$ is feasible, similarly, the residual $R=I-D_{\sigma}$ can be FNE in (7) with $\mathcal{M}=I$.
@@
 3 Halpern-type preconditioned proximal point (HPPP)
-Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29].
-The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [26] is an implicitly regularized method for finding a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$.
+Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [he2024convergence].
+The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [halpern1967fixed] is an implicitly regularized method for finding a particular fixed point [Diakonikolas2020], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [halpern1967fixed, lions1977approximation, Wittmann1992, 2002Iterative, xu2004viscosity, Qi2021, he2024convergence], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$.
 When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$.
@@
 Based on the definition of the upper limit, we take a subsequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ such that
-(12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
+[…]
 By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, and we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$.
@@
 It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfies conditions (i)-(iii) of Theorem 3.1.
-Compared with [3], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point.
-The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$.
+Compared with [Bauschke2023], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point.
+The work [Bauschke2023] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$.
 By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
@@
 $\mathcal{A}\mathcal{T}\mathbf{u}^{k_{n}}\owns\mathcal{M}(\mathbf{u}^{k_{n}}-\mathcal{T}\mathbf{u}^{k_{n}})\to\mathbf{0}.$
-Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$.
+Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [BCombettes, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$.
 Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$.
@@
 Remark 3.5.
-We compare Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
+We compare Theorem 3.1, Proposition 3.3 with [Bredies2022, Theorem 2.9] and [Bredies2022, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$.
 All conditions are the same, except for the additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$.
 The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption, especially in applications to splitting algorithms, and is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
-Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [31], where $U$ is a closed convex cone and $V$ is a closed hyperplane.
-Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence.
+Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [HUNDAL200435], where $U$ is a closed convex cone and $V$ is a closed hyperplane.
+Building on Hundal’s counterexample, [DR_weak_convergence2020] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence.
 Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
@@
 Proof 3.6.
-As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
+As proved in [DR_weak_convergence2020], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
 (15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k},$
@@
 and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$.
-By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$.
+By [Bredies2022, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$.
 Indeed, if by contraction, $\mathbf{u}^{k}\to\mathbf{0}$, then
@@
 Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2).
-The restarted technique is an effective method in machine learning [41] and LP [17, 38].
+The restarted technique is an effective method in machine learning [Park2022] and LP [chen2024hprlpimplementationhprmethod, lu2024restartedhalpernpdhglinear].
 Algorithm 2 Restarted HPPP
@@
 7:end for
-8:$\mathbf{u}^{N_{e},q}$
+[…]
 3.2 Convergence rate
-Based on the following technical lemma, Sabach and Shtern [50] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
-Lemma 3.9 ([50]).
+Based on the following technical lemma, Sabach and Shtern [Sabach2017] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
+Lemma 3.9 ([Sabach2017]).
 Let $M_{1}>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers such that $a_{1}<M_{1}$ and
@@
 $\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}\left(\frac{1}{k}\right).$
-Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3].
-By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE.
+Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [Bredies2022, Proposition 2.3].
+By [Bredies2022, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE.
 Let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, and then HPPP is equivalent to the following reduced algorithm:
 (16) $\mathbf{w}^{0}=\mathcal{C}^{*}\mathbf{u}^{0},\mathbf{w}^{k+1}=\mu_{k+1}\mathbf{w}^{0}+(1-\mu_{k+1})\tilde{\mathcal{T}}\mathbf{w}^{k}.$
-By [36, Theorem 2.1] and [52, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
+By [lieder2021convergence, Theorem 2.1] and [Sun2024, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
 Proposition 3.12.
@@
   1. (i)
-    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$,
+[…]
   2. (ii)
@@
 Under the ideal conditions of RED and with assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$.
-Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17).
+Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [beck2017first, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17).
 Set $\mathbf{K}=I$; the implicit gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3.
-Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers.
+Even if the conditions of RED are difficult to meet [REDClarifications2019], recent developments [ryu2019plug, building_FNE, MMO2021, pmlr-v202-delattre23a, BUI2020124315, bredies2024learningfirmlynonexpansiveoperators, Wei2024] can learn FNE denoisers.
 According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
 3.3.1 Acceleration for PnP methods
-PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$.
+PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [ryu2019plug, hurault2022proximal], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$.
 When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration.
-Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9].
+Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [Bredies2015, Bredies2022].
 The related GraRED-P3 algorithm is shown in Algorithm 4.
@@
 Remark 3.13.
-In [28], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM.
+In [He2015], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM.
 In contrast, GraRED-HP3 obtains an accelerated convergence rate of $\mathcal{O}\left(\frac{1}{k}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
@@
 6:end for
-7:$\mathbf{x}^{N}$.
+[…]
 Algorithm 4 GraRED-P3
@@
 6:end for
-7:$\mathbf{x}^{N}$.
+[…]
 Algorithm 5 Relaxed RED-PRO via SD
@@
 3.3.2 Compared with RED-PRO
-RED-PRO [20] is defined by
+RED-PRO [cohen2021regularization] is defined by
 (18) $\min_{\mathbf{x}\in\mathrm{Fix}(T)}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}.$
@@
 where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, and we plot $f(x)+g(x)$ in Figure 1.
-[…]
+2-21-111221122f(x)+g(x)f(x)+g(x)xxyy
 Figure 1: The image of $f(x)+g(x)$.
-The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$.
+The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval [1,0][-1,0].
 We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$.
@@
 $\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0,\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right$.
-If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction.
+If $x^{*}=1$, assume that $1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$, and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to a contradiction.
 Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
@@
 (21) $P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0,\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
-$x-y-1\leq 0$$x$$y$$O$$S(2,0)$$R(3,1)$
+$1\leq 0xxyy$OOS(2,0)S(2,0)$R(3,1)$
 Figure 2: $\mathcal{M}$-projection onto the saddle-point set.
@@
 4.2 A 2D toy example
-We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
+We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [Park2022] for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
 $T_{\theta}\mathbf{x}=\frac{1}{\gamma}\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\mathbf{x}.$
@@
 4.3 An infinite-dimensional example
-We consider an infinite-dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences.
+We consider an infinite-dimensional problem on $\ell^{2}$ from [bc24], the space of square summable sequences.
 For any $\mathbf{x}=(x_{0},x_{1},x_{2},\ldots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$.
@@
 where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\ldots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\ldots)$.
-Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29].
+Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [BCombettes, Example 20.29].
 Additionally, by solving the above inclusion problem (22), we obtain $\mathbf{x}^{*}=(1,0,\ldots)$.
@@
 4.4 Image deblurring
-First, we compared CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
+First, we compared CP [Chambolle2011], PPP [Bredies2022], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
 (23) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|_{2}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
@@
 Refer to captionRefer to caption
-(c) CP ($27.85$dB)
+(c) CP (27.8527.85dB)
 Refer to captionRefer to caption
-(d) $\lambda_{k}=1.95$: PPP ($27.85$dB)
+(d) $\lambda_{k}=1.95$: PPP (27.8527.85dB)
 Refer to captionRefer to caption
-(e) $\lambda_{k}=1.2$: PPP ($27.85$dB)
+(e) $\lambda_{k}=1.2$: PPP (27.8527.85dB)
 Refer to captionRefer to caption
-(f) HPPP ($27.99$dB)
+(f) HPPP (27.9927.99dB)
 Refer to captionRefer to caption
-(g) GraRED-P3 ($29.70$dB)
+(g) GraRED-P3 (29.7029.70dB)
 Refer to captionRefer to caption
-(h) GraRED-HP3 ($29.67$dB)
+(h) GraRED-HP3 (29.6729.67dB)
 Figure 6: Deblurring results of Butterfly degraded by the Gaussian blur kernel with noise level $\sigma=2.55$.
@@
 Second, we compared RED and RED-PRO.
-Following [47, 20], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
+Following [Romano2017, cohen2021regularization], a $9\times 9$ uniform point spread function (PSF) and a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images.
 We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$.
@@
 Figure 8: Recovery performance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
-Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO.
-We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[65] and GS ($D_{3}$[32, 33].
-The PnP-DYS method [59] uses the Davis-Yin splitting (DYS) algorithm.
+Additionally, we trained a FNE-DnCNN ($D_{1}$) using real spectral normalization [ryu2019plug], which is employed in RED and RED-PRO.
+We also compared our results with state-of-the-art methods using advanced denoisers, such as DRUNet ($D_{2}$[zhang2021plug] and GS ($D_{3}$[hurault2022gradient, hurault2022proximal].
+The PnP-DYS method [There-Operator-Splitting2024] uses the Davis-Yin splitting (DYS) algorithm.
 For GraRED-HP3, the anchor point is the same as the initial point.
@@
 The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
-Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [18]. The best two results are highlighted in red and blue, respectively.
+Table 3: Recovery results are obtained by different algorithms with trainable nonlinear reaction-diffusion (TNRD) denoiser [Chen2017].
+The best two results are highlighted in red and blue, respectively.
 Algorithms Uniform kernel Gaussian kernel($\sigma_{k}=1.6$)
 Bike Butterfly Flower Hat Bike Butterfly Flower Hat
-RED [47] 26.10 30.41 30.18 32.16 27.90 31.66 32.05 33.30
-RED-PRO(HSD) [20] 24.95 27.24 29.38 31.55 27.36 30.55 31.81 33.07
-RRP [20] 26.48 30.64 30.46 32.25 28.02 31.66 32.08 33.26
+RED [Romano2017] 26.10 30.41 30.18 32.16 27.90 31.66 32.05 33.30
+RED-PRO(HSD) [cohen2021regularization] 24.95 27.24 29.38 31.55 27.36 30.55 31.81 33.07
+RRP [cohen2021regularization] 26.48 30.64 30.46 32.25 28.02 31.66 32.08 33.26
 GraRED-P3 26.55 30.72 30.67 32.43 28.13 31.81 32.42 33.51
@@
 […]
-Restarted HPPP$(D_{2})$ $\tau=1.2,s=1/1.2,N=30,q=10$
+[…]
 […]
@@
 Pepper Craft Cameraman Couple Man House Starfish Butterfly Boat
-RED [47] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
-RED-PRO [20] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
-DPIR [65] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
-PnP-DRS [33] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
-PnP-PGD [32] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
-PnP-DYS [59] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
+RED [Romano2017] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
+RED-PRO [cohen2021regularization] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
+DPIR [zhang2021plug] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
+PnP-DRS [hurault2022proximal] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
+PnP-PGD [hurault2022gradient] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
+PnP-DYS [There-Operator-Splitting2024] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
 […]
-Restarted HPPP$(D_{2})$ 28.34 26.75 27.02 29.71 30.44 32.69 28.91 29.28 30.01
 […]
-Restarted HPPP$(D_{3})$ 29.81 27.13 27.49 30.04 30.80 33.15 30.09 30.53 30.37
+[…]
+[…]
 Figure 9: Deblurring results of Butterfly degraded by the uniform kernel.
 4.5 Image inpainting
-In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [16], PPP [9], and HPPP algorithms.
+In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [Chambolle2011], PPP [Bredies2022], and HPPP algorithms.
 The discrete image inpainting model is
@@
   • CP: $\tau=s=1/\left\|K\right\|=0.57$.
-Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used:
+Both GraRED-P3 and GraRED-HP3 use DnCNN [ryu2019plug], other parameters are used:
   • GraRED-P3 : $\tau=10,s=0.1,\lambda_{k}=0.2,\lambda=5$;
@@
 Refer to captionRefer to caption
-(c) CP ($28.94$dB)
+(c) CP (28.9428.94dB)
 Refer to captionRefer to caption
-(d) $\lambda_{k}=1.6$: PPP ($29.06$dB)
+(d) $\lambda_{k}=1.6$: PPP (29.0629.06dB)
 Refer to captionRefer to caption
-(e) $\lambda_{k}=1.2$: PPP ($29.09$dB)
+(e) $\lambda_{k}=1.2$: PPP (29.0929.09dB)
 Refer to captionRefer to caption
-(f) HPPP ($29.20$dB)
+(f) HPPP (29.2029.20dB)
 Refer to captionRefer to caption
-(g) GraRED-P3 ($36.15$dB)
+(g) GraRED-P3 (36.1536.15dB)
 Refer to captionRefer to caption
-(h) GraRED-HP3 ($36.14$dB)
+(h) GraRED-HP3 (36.1436.14dB)
 Figure 10: Recovery results of House degraded by the random mask.
@@
 Refer to captionRefer to caption
-(k) CP ($31.58$dB)
+(k) CP (31.5831.58dB)
 Refer to captionRefer to caption
-(l) $\lambda_{k}=1.6$: PPP ($31.79$dB)
+(l) $\lambda_{k}=1.6$: PPP (31.7931.79dB)
 Refer to captionRefer to caption
-(m) $\lambda_{k}=1.2$: PPP ($31.74$dB)
+(m) $\lambda_{k}=1.2$: PPP (31.7431.74dB)
 Refer to captionRefer to caption
-(n) HPPP ($32.00$dB)
+(n) HPPP (32.0032.00dB)
 Refer to captionRefer to caption
-(o) GraRED-P3 ($37.46$dB)
+(o) GraRED-P3 (37.4637.46dB)
 Refer to captionRefer to caption
-(p) GraRED-HP3 ($37.47$dB)
+(p) GraRED-HP3 (37.4737.47dB)
 Figure 11: Recovery results of House degraded by the character mask.
@@
 Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments.
-In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [4].
+In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [Bauschke2021].
 Acknowledgments
 The authors thank Prof.
-Defeng Sun for bringing the interesting work [52] to our attention and Mr.
+Defeng Sun for bringing the interesting work [Sun2024] to our attention and Mr.
 Feng Xue for his insightful discussions.
@@
 First, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5).
-To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [30, Lemma 2.5].
+To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce an important lemma from [2002Iterative, Lemma 2.5].
 Lemma A.1.
@@
 Proof A.3.
-Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [9, Proposition 2.3]. Denote
+Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [Bredies2022, Proposition 2.3]. Denote
 (25) $C=L\left\|\mathcal{C}\right\|,$
@@
 where the third inequality follows from the $L$-Lipschitz property of $(\mathcal{M}+\mathcal{A})^{-1}$, and the fourth equality follows from the definition of $\left\|\mathbf{u}\right\|_{\mathcal{M}}$, i.e., $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}=\left\|\mathcal{C}^{*}\mathbf{u}\right\|$ for any $\mathbf{u}\in\mathcal{H}$.
-First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$, we have
+First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, we have
 $\left\lVert\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\rVert_{\mathcal{M}}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}$
@@
 It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$, which completes the proof of Lemma B.1.
-As mentioned in [3, 15], we can introduce the following notion of $\mathcal{M}$-projection.
+As mentioned in [Bauschke2023, BUI2020124315], we can introduce the following notion of $\mathcal{M}$-projection.
 Definition B.3 ($\mathcal{M}$-projection).